### v1

# Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems

Abstract
Recent advancements in Large Language Model (LLM)-based Multi-Agent Systems (MAS) have demonstrated remarkable potential for tackling complex decision-making tasks. However, existing frameworks inevitably rely on serialized execution paradigms, where agents must complete sequential LLM planning before taking action. This fundamental constraint severely limits real-time responsiveness and adaptation, which is crucial in dynamic environments with ever-changing scenarios. In this paper, we propose a novel parallelized planning-acting framework for LLM-based MAS, featuring a dual-thread architecture with interruptible execution to enable concurrent planning and acting. Specifically, our framework comprises two core threads: (1) a planning thread driven by a centralized memory system, maintaining synchronization of environmental states and agent communication to support dynamic decision-making; and (2) an acting thread equipped with a comprehensive skill library, enabling automated task execution through recursive decomposition. Extensive experiments on challenging Minecraft demonstrate the effectiveness of the proposed framework.
Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems
[…]
1 Introduction
Figure 1: An overview of the proposed multi-agent system
Multi-Agent Systems (MAS) have become a well-established paradigm for tackling complex decision-making problems Hong et al. (2023); Chen et al. (2024b); Dong et al. (2024), with early efforts primarily relying on reinforcement learning Busoniu et al. (2008); Yang and Wang (2021); Lowe et al. (2017) to enable multiple agents to cooperate or compete in dynamic environments. Despite the encouraging results achieved, these MAS frameworks faced limitations in handling complex real-world scenarios that require advanced communication, reasoning, and adaptability. The rapid advancement of Large Language Models (LLMs) DeepSeek-AI (2025, 2024); Brown et al. (2020); Achiam et al. (2023); Qwen et al. (2024); Dubey et al. (2024) has since revolutionized MAS by adding natural language understanding and generation capabilities, enabling agents to engage in more sophisticated collaboration. LLMs have significantly enhanced the flexibility and versatility of MAS, opening the door to more complex tasks and dynamic interactions in real-world applications Wu et al. (2024); D’Arcy et al. (2024); Chen et al. (2024c); AL et al. (2024).
Recent works have demonstrated the potential of LLM-based MAS in various domains. MetaGPT Hong et al. (2023) introduces an innovative meta-programming framework to enhance task decomposition and agent collaboration, AgentVerse Chen et al. (2024b) improves collaborative performance by orchestrating expert agents, and VillagerAgent Dong et al. (2024) tackles task dependencies in complex environments using DAG-based task decomposition. Despite these advancements, most current frameworks still rely on serialized execution, where planning and acting occur sequentially for each agent. This serialized nature creates a substantial bottleneck when handling dynamic information, particularly evident in dynamic settings like Minecraft, where the environment is constantly changing, even without the intervention of agents. Although Voyager Wang et al. (2023) pioneered LLMs-based agents in Minecraft, its approach, which pauses the game server during LLM interactions, effectively staticizes the dynamic environment, thus failing to fully address the real-time responsiveness needed in such settings. Subsequent works in Minecraft agents Liu et al. (2024); Zhao et al. (2024) have followed this approach, but they also inherit the same limitations.
Our analysis reveals three critical challenges in current LLM-based MAS for dynamic environments. First, inflexible action scheduling is prevalent, as many existing agent frameworks rely on serialized execution, requiring agents to wait for a language model response before proceeding with further actions. This rigidity complicates the handling of unexpected environmental changes; Second, limited replanning capabilities hinder agents’ performance, as they often execute actions to completion without interruption. This lack of adaptability prevents agents from effectively reconsidering or adjusting their plans in response to urgent and unforeseen events, diminishing their overall effectiveness. Lastly, memory sharing delays pose another issue, as memory updates in many multi-agent systems only occur after an action has been fully executed. This results in delayed observational data sharing, causing agents to operate based on outdated information, which in turn limits the team’s coordination and efficiency.
In this paper, we propose a parallelized planning-acting framework that introduces a dual-thread architecture with interruptible execution for efficient LLM-based MAS in dynamic environments, as shown in Fig. 1. Our architecture decouples LLM reasoning from action execution, enabling concurrent planning and acting. Moreover, the interruption mechanism enables agents to adjust their actions in real time based on environmental changes, thereby improving their adaptability. Specifically, our framework consists of two core threads: (1) A planning thread employing a centralized memory system to support efficient and timely information sharing among agents, minimizing memory sharing delays and ensuring agents operate with up-to-date information for better coordination and efficiency. (2) An acting thread utilizing a comprehensive skill library, enabling efficient task execution through a recursive task decomposition mechanism. Our core contributions are summarized as follows:
  • We propose a parallelized planning-acting framework that decouples planning and acting into parallelized dual threads with interruptible execution for efficient LLM-based MAS.
  • We develop a centralized memory system to support the planning thread, ensuring that agent decisions are always informed by the latest environmental changes and interactions.
  • We design a comprehensive skill library to empower the acting thread, enabling efficient task execution through recursive task decomposition.
  • Experimental results on Minecraft demonstrate a paradigm shift from serialized deliberation to parallelized interaction, yielding notable improvements in efficiency and coordination.
2 LLM-based Multi-Agent Framework
We propose a novel parallelized planning-acting framework based on LLMs, designed to enable collaboration and decision-making among agents in complex dynamic environments. Our framework introduces three key innovations: (1) A dual-thread architecture with an interruptible execution mechanism, enabling concurrent planning and acting, (2) A real-time updated centralized memory system supporting the planning thread, ensuring that agents’ decisions are informed by the latest environmental changes and team interactions, and (3) A Minecraft comprehensive skill library supporting the acting thread, automating task execution by proposing a recursive task decomposition mechanism. While the skill library represents an engineering contribution, it is important to note that our framework is generic and can be extended to other domains with minimal adaptation. This flexibility highlights the broader applicability of our approach, while Minecraft provides a rich and accessible testbed for evaluating its performance in dynamic environments.
2.1 Parallelized Planning-Acting Framework
Inspired by the human ability to think and act simultaneously, our framework features a dual-thread architecture in Fig. 2 that separates planning (supported by LLMs and the centralized memory system) and acting (executed by the comprehensive skill library). Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{n}\}$ denote the set of agents. Our framework operates as follows:
[…]
    $A_{i}^{t+1}=\text{LLM}(S,O_{i}^{t},C^{t},A_{i}^{t}).$ (1)
    The planned action $A_{i}^{t+1}$ is then written into an action buffer, which serves as the communication channel between the planning thread and the acting thread. This action buffer is implemented as a single-slot queue, allowing the planning thread to write and the acting thread to read. If the buffer is already occupied, the previous action is discarded to make space for the new one. This mechanism ensures that the planning thread always places the most up-to-date action (based on the latest observations and agent status) into the buffer, while the acting thread always retrieves the most current action, never operating outdated plans.
  • Acting Thread: The acting thread retrieves actions from the action buffer and executes them according to the following rule:
    $s_{i}=\begin{cases}A_{i}^{t+1}&\text{if }p(A_{i}^{t+1})>p(A_{i}^{t}),\\ A_{i}^{t}&\text{otherwise}.\end{cases}$ (2)
    Here, $s_{i}$ represents the skill execution of agent $a_{i}$, $A_{i}^{t}$ is the current action being executed, $A_{i}^{t+1}$ is the new action in the buffer, and $p$ is the priority of an action. If an interruption is triggered—i.e., when the LLM determines that the new action has a higher priority than the current action—the planning thread sends an interrupt signal to the acting thread. The acting thread immediately exits and restarts with the new action. If no interruption is triggered, the acting thread completes the current action $A_{i}^{t}$ before retrieving and executing the next action $A_{i}^{t+1}$ from the buffer.
Latency Analysis. The parallelized architecture intuitively reduces system latency through concurrent execution of planning and acting threads. Let $T_{\text{plan}}$ denote the LLM reasoning latency and $T_{\text{act}}$ the skill execution time. For a task requiring $n$ atomic actions without any interruption:
  • Serialized Framework:
    $T_{\text{s}}=\sum_{k=1}^{n}(T_{\text{plan}}^{(k)}+T_{\text{act}}^{(k)}).$ (3)
  • Parallelized Framework:
    $T_{\text{p}}=T_{\text{plan}}^{(1)}+\sum_{k=2}^{n}\max(T_{\text{plan}}^{(k)},T_{\text{act}}^{(k-1)})+T_{\text{act}}^{(n)}.$ (4)
The latency reduction $\Delta T$ can be expressed as:
$\Delta T\approx\sum_{k=1}^{n}\left(T_{\text{plan}}^{(k)}+T_{\text{act}}^{(k)}-\max(T_{\text{plan}}^{(k)},T_{\text{act}}^{(k)})\right).$ (5)
This analysis highlights two key advantages of our framework: 1) The initial planning latency $T_{\text{plan}}^{(1)}$ is effectively amortized over subsequent actions, and 2) The overlapping of planning and acting phases successfully conceals $T_{\text{plan}}$ when $T_{\text{act}}>T_{\text{plan}}$ (our proposed comprehensive skill library in Section 2.3 ensures this condition is well-maintained). Our experimental results in Section 3.4 demonstrate the overall efficiency advantage of the parallelized framework.
By decoupling planning and acting into parallelized threads, our framework strikes a balance between efficiency and flexibility, allowing agents to respond dynamically to unpredictable environments with enhanced adaptability and performance.
2.2 Centralized Memory System
Figure 2: An overview of the proposed parallelized planning-acting architecture
To facilitate effective coordination, we implement a centralized memory system $M$ that stores and manages information at the team level. The memory is updated at each time step $t$ as follows:
$M^{t+1}=\{O^{t+1},C^{t+1},A^{t+1}\}\cup\{M^{t}\setminus O^{t}\},$ (6)
where $O^{t+1}$ denotes the updated observations of the multi-agent system at time $t+1$, which overwrite the previous observations $O^{t}$, $C^{t}$ denotes the chat messages of the system at time $t$, $A^{t}$ denotes the action history of the system at time $t$. This unified repository enables agents to access and utilize relevant information during task execution, ensuring efficient team coordination:
  • Observation Records: Each agent’s observations are continuously updated in the centralized memory, reflecting the latest agent status and environmental state. These observations are associated with the respective agent, allowing the team to maintain a comprehensive and up-to-date view of the environment.
  • Chat Logs: All team chat messages are stored in the centralized memory, with long-term retention to support historical analysis and decision making. During planning, agents can retrieve the most recent chat messages to incorporate team insights into their strategies. This ensures that decisions are informed by the collective knowledge of the team.
  • Action History: The centralized memory also records the actions taken by each agent, providing a detailed history of task execution. These logs are useful for behaviours refining and performance analysis.
We implement two types of multi-agent communication (passive and active) ensuring that the centralized memory remains a dynamic resource for team coordination:
  • Passive Communication: In the planning thread, after each planning cycle, the LLM generates a chat message based on the agent’s latest observations, which is then sent to the centralized memory’s chat logs. This ensures that passive communication, reflecting updated observations, can run concurrently with action execution. While an agent is performing actions, its observations are continuously updated and shared with the team, enabling real-time coordination based on the most current environmental information.
  • Active Communication: In the acting thread, agents can actively choose to send chat messages by performing a chat action. This action allows the agent to share any information with teammates, updating the chat logs in real-time. This form of communication ensures that agents can respond dynamically and share critical information during action execution, facilitating efficient and up-to-date information exchange between agents.
2.3 Comprehensive Skill Library
To enable seamless interaction between agents and the Minecraft environment, we develop a comprehensive skill library that encapsulates a wide range of in-game actions. The library provides high-level APIs for tasks such as resource collection, combat, exploration, and communication. For further technical details, please refer to Appendix B.
A key feature of the comprehensive skill library is the implementation of a recursive task decomposition mechanism, which automates the completion of prerequisite tasks such as mining raw materials and crafting necessary tools. This automation ensures that agents can perform complex resource collection tasks with minimal manual intervention, enabling the automated collection of over 790 types of items in Minecraft, surpassing all existing methods  Wang et al. (2023); Zhu et al. (2023); Zhao et al. (2024); Liu et al. (2024).
The core recursive process can be formally modeled as a weighted directed acyclic graph (DAG) $\mathcal{G}=(V,E,\phi)$, where:
  • Vertex set $V=\{v_{i}\}$ represents atomic tasks:
    $v_{i}=(t_{i},c_{i},f_{i}),$ (7)
    where $t_{i}\in\mathcal{T}$ is the target item type (All collectible items in Minecraft), $c_{i}\in\mathbb{N}^{+}$ denotes required quantity, and $f_{i}$ specifies the operation type for obtaining the item.
  • Edge set $E\subseteq V\times V$ encodes prerequisite task dependencies:
    $(v_{j},v_{i})\in E\iff v_{j}\in\text{pre}(v_{i}),$ (8)
    where $\text{pre}(v_{i})$ gives the prerequisite tasks for producing $t_{i}$.
  • Weight function $\phi:E\to\mathbb{Q}^{+}$ defines material conversion rates:
    $\phi(v_{j},v_{i})=\frac{r_{ij}}{n_{\text{out}}},$ (9)
    with $r_{ij}$ being the required quantity of $t_{j}$ per operation, and $n_{\text{out}}$ being the output quantity per operation.
The recursive resolution process follows:
$\Psi(v_{i})=\bigcup_{(v_{j},v_{i})\in E}\left\{\Psi\left(v_{j}^{\left(\phi(v_{j},v_{i})\cdot c_{i}\right)}\right)\right\}\cup\{v_{i}\},$ (10)
where $v_{j}^{(k)}$ denotes a task requiring $k$ units of $t_{j}$, and the base case $\Psi(v_{i})=\emptyset$ applies when $I(t_{i})\geq c_{i}$, with $I$ representing the current inventory state, which means the task is accomplished.
3 Experiments
3.1 Experiment Setup
We aim to address the following questions through further experimentation: (1) Can our comprehensive skill library enable agents to efficiently accomplish basic tasks within Minecraft’s open-world environment? (2) How well can our proposed multi-agent framework perform on highly challenging and complex tasks? (3) Can our proposed parallelized planning-acting framework better support multi-agent collaboration, competition, and human-agent interaction?
All experiments are conducted in the Minecraft Java Edition version 1.19.4 gaming environment using the Qwen-Plus model  Qwen et al. (2024), while multi-modal experiments utilized the Qwen-VL-Plus model  Bai et al. (2023). The game server operates continuously without pausing during interactions with LLMs, thereby necessitating that all agents perform in real-time. All evaluation metrics are captured dynamically during these interactions. Metrics such as ’time’ denote the total duration required by the multi-agent systems to complete their designated tasks, encompassing the time expended on communications with LLMs.
3.2 Benchmark Task Scenarios
To evaluate the generalizability of our multi-agent framework in Minecraft, we designed three distinct task scenarios: Resource Collection, Boss Combat, and Adversrial Player versus Player (PvP). These scenarios were selected to cover a range of application domains, showcasing the versatility of our multi-agent framework across diverse contexts.
  • Resource Collection: This task involves gathering resources in the Minecraft world, evaluating the efficiency of the comprehensive skill library and the coordination of multi-agent system in performing repetitive and structured tasks.
  • Boss Combat: This task requires the multi-agent system to defeat a strong boss-level enemy in a predefined, highly intricate combat environment, evaluating advanced strategic planning, real-time decision-making, effective collaboration among agents and the system’s ability to handle dynamic and challenging environments.
  • Adversarial PVP: This task focuses on team-based adversarial competition, where two teams of agents compete against each other in a combat scenario. The primary objective is to evaluate the agents’ ability to strategize, adapt, and outperform opponents.
3.3 Main Results
3.3.1 Resource Collection Task
We first evaluate the effectiveness of our skill library’s recursive task decomposition mechanism by assessing a single agent’s performance on basic resource collection tasks in Minecraft. This initial evaluation aimed to verify the comprehensive skill library’s capability in automating complex workflows, thereby validating its efficiency and reliability in handling fundamental Minecraft tasks. As shown in Table 1, when the recursive task decomposition mechanism is ablated, the system can only complete short-term tasks that require fewer steps and shows a reduction in efficiency. In contrast, utilizing our comprehensive skill library enables efficient completion of all tasks.
Table 1: Completion time and success rate (SR) comparison with and without the Recursive Task Decomposition Mechanism (RTDM). All results are shown with the mean and standard deviation over 10 trials.
Task with RTDM w/o RTDM
Time(min) SR Time(min) SR
[Uncaptioned image] $0.3\pm 0.2$ $100\%$ $2.8\pm 2.5$ $100\%$
[Uncaptioned image] $0.6\pm 0.3$ $100\%$ $4.6\pm 2.5$ $80\%$
[Uncaptioned image] $1.4\pm 0.5$ $100\%$ $N/A$ $0\%$
[Uncaptioned image] $4.7\pm 1.3$ $100\%$ $N/A$ $0\%$
[Uncaptioned image] $6.2\pm 1.6$ $100\%$ $N/A$ $0\%$
We then evaluate our framework’s performance on eight more complicated resource collection tasks. Table 2 compares the completion times between our multi-agent system (3 agents) and single-agent baseline, demonstrating the efficiency gains through coordination.
Table 2: Completion time (in minutes) of resource collection tasks comparison between multi-agent(MA) and single-agent(SA) systems. All results are shown with the mean and standard deviation over 10 trials. For detailed task definitions, please refer to Appendix C.
Task MA Time SA Time
Iron Tool Set [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] $7.8\pm 2.1$ $8.5\pm 3.7$
Diamond Armor [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $13.7\pm 4.1$ $28.3\pm 6.1$
Redstone Devices [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $11.0\pm 6.0$ $13.1\pm 3.3$
Navigation Kit [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $25.3\pm 12.2$ $39.4\pm 11.7$
Transport System [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $22.0\pm 10.1$ $37.8\pm 12.6$
Food Supplies [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $6.6\pm 3.9$ $8.0\pm 2.0$
Building Materials [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $15.8\pm 2.8$ $22.6\pm 7.4$
Storage System [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $10.0\pm 8.9$ $16.7\pm 7.8$
The experimental results demonstrate that our Minecraft comprehensive skill library, when employed within a multi-agent framework, can efficiently complete various resource collection tasks as well. As shown in the comparisons, the system with three agents significantly reduces task completion times compared to the single-agent baseline across most tasks, validating the effectiveness of our approach and showcasing the efficiency benefits of multi-agent collaboration. However, the observed non-linear scaling (multi-agent time > single-agent time/3) does not always achieve ideal linear speedup due to two inherent constraints:
  • Resource dependency chains: Sequential prerequisites in crafting workflows (e.g., iron tools required for diamond mining)
  • Spatial contention: Overlapping access to resources in close proximity (e.g., multiple agents competing to gather items from the same ore cluster)
3.3.2 Boss Combat Task
We conduct comprehensive evaluations across three challenging combat scenarios with varying agent team sizes. Experiments are conducted in three predefined scenarios, each featuring an extremely powerful and representative boss from Minecraft’s major dimensions: the Elder Guardian (Overworld), the Wither (Nether), and the Ender Dragon (End). The initial inventory of agents remained consistent across all scenarios. Their equipment are validated to be collectible within a short time during the resource collection task, and other consumable resources can also be efficiently gathered ultilizing our skill library. For further details, please refer to Appendix C.
The performance of the multi-agent system in the boss combat task is summarized in Table 3, demonstrating that our framework achieves high success rates in completing all challenging boss combat tasks with various agent team sizes.
Table 3: Boss combat performance across three task scenarios. All results are shown with the mean and standard deviation over 12 trials. For detailed evaluation metrics definitions, please refer to Appendix C.
Task Scenario # Agents Time(min) Health Ratio Progress SR
Elder Guardian (Overworld) [Uncaptioned image] 3 $2.4\pm 2.1$ $49.8\pm 32.8\%$ $91.4\pm 20.1\%$ $83.3\%$
5 $1.2\pm 0.8$ $84.4\pm 7.9\%$ $100.0\pm 0.0\%$ $100.0\%$
10 $1.2\pm 0.3$ $86.9\pm 8.4\%$ $100.0\pm 0.0\%$ $100.0\%$
Wither (the Nether) [Uncaptioned image] 3 $1.8\pm 0.8$ $18.6\pm 23.4\%$ $71.8\pm 35.9\%$ $41.7\%$
5 $1.5\pm 0.5$ $53.4\pm 32.7\%$ $88.1\pm 20.6\%$ $75.0\%$
10 $1.4\pm 0.3$ $69.5\pm 19.0\%$ $100.0\pm 0.0\%$ $100.0\%$
Ender Dragon (the End) [Uncaptioned image] 3 $N/A$ $0.0\pm 0.0\%$ $20.4\pm 18.5\%$ $0.0\%$
5 $6.5\pm 2.0$ $18.1\pm 23.2\%$ $75.4\pm 28.9\%$ $41.7\%$
10 $5.2\pm 1.5$ $43.9\pm 24.4\%$ $98.2\pm 5.9\%$ $91.7\%$
3.3.3 Adversarial PVP Task
In the adversarial player-vs-player (PVP) task, we conduct a direct comparison between the parallelized and serialized frameworks across various team sizes. As shown in Fig. 3, this setup clearly demonstrate and quantify the differences of these two methodologies under competitive conditions.
As shown in Fig. 3, the parallelized framework demonstrates a significant advantage over the serialized framework in this dynamic adversarial scenario. Our analysis indicates that this advantage is primarily attributed to our interruption mechanism, which enables agents to dynamically adjust their strategies and respond promptly to changes in the environment (e.g., seamlessly switch attack targets, prioritize health restoration, etc.).
Figure 3: Comparison of parallelized vs. serialized frameworks in PVP tasks. All results are shown with the mean and standard deviation over 10 trials.
3.4 Ablation Study
To validate the necessity of our method, we perform ablation studies in boss combat tasks by disabling different components in our framework:
  • w/o Parallelized Planning-Acting Framework: Replaced our concurrent architecture with traditional serialized execution (LLM call → action execution), disabling action interruption capabilities once execution begins.
  • w/o Centralized Memory System: Disabled real-time team observation polling, team chat logs, action history and global progress information, thereby restricting agents to rely solely on their individual observations.
  • Multimodal Observation: Replaced text-based observations with visual inputs using Mineflayer’s prismarine-viewer, capturing first-person screenshots processed by a vision-language model (VLM) as observation.
These ablation configurations systematically evaluate the contribution of each architectural component to overall system performance, providing insights into their relative importance in complex multi-agent coordination tasks.
Experiment results shown in Fig. 4 highlight the critical roles of the parallelized planning-acting framework and centralized memory system in maintaining this high performance. When we replaced the text-based observation space with a multimodal observation system, we observed that although the inclusion of visual language models (VLMs) increased response latency and lowered the observation accuracy, the agents still performed reasonably well. This indicates that our framework remains robust even when handling different modalities.
Figure 4: Ablation study in boss combat tasks across 12 trials in three scenarios: Elder Guardian (EG), Wither (Wi), and Ender Dragon (ED), where efficiency is defined as the inverse of completion time in minutes.
3.5 Human-Agent Interaction
Our framework also supports flexible human-agent interaction, allowing human players to either actively participate in task execution alongside agents, facilitating dynamic cooperation within the game environment, or to take on a guiding role by providing instructions to agents for task division, coordination, and strategic decision-making. Further experimental cases demonstrating the effectiveness of human-agent interaction are presented in Appendix C.4.
4 Related Works
LLM-based Minecraft Agents. The development of LLM-based AI agents in Minecraft has evolved through several key approaches: Voyager  Wang et al. (2023) established the first LLM-based agent with automatic skill discovery using GPT-4  Achiam et al. (2023). Subsequent studies enhanced agents via specialized memory mechanisms  Park et al. (2024); Li et al. (2024a), specialized LLM fine-tuning  Feng et al. (2023); Zhao et al. (2024); Liu et al. (2024), task decomposition and causal graph learning  Yuan et al. (2023); Zhu et al. (2023); Yu and Lu (2024), and combination with reinforcement learning  Li et al. (2024b, 2023). Additionally, multi-modal information perception and processing were explored  Zheng et al. (2023); Cai et al. (2024), along with other novel techniques  Wang et al. (2024); Zhou et al. (2024). The development of benchmarks for general capabilities progressed with MineDojo  Fan et al. (2022) and MCU  Lin et al. (2023), while specific agent capabilities were assessed through additional benchmarks  Qin et al. (2024); Dong et al. (2024).
LLM-based Multi-Agent Systems. Recent surveys  Mou et al. (2024); Guo et al. (2024) provide a comprehensive overview of the key advancements in LLM-based multi-agent systems. Building on these foundations, recent research has focused on several core areas: infrastructure frameworks for efficient agent coordination  Gong et al. (2024); Chen et al. (2024b); Zhang et al. (2024b); Hong et al. (2023); Zhang et al. (2024a), which introduce novel paradigms for task management and team collaboration; benchmark development to evaluate multi-agent performance in dynamic environments  Chen et al. (2024a); Dong et al. (2024), which has created robust testing environments to assess the generalization and efficiency of LLM-powered agents; large-scale social simulations  AL et al. (2024); Park et al. (2023); Yang et al. (2024), which explore how multi-agent systems model complex societal behaviors; and domain-specific applications  Wu et al. (2024); D’Arcy et al. (2024); Chen et al. (2024c) demonstrating the effectiveness of LLM-based agents in specific scenario simulation.
5 Conclusion
We propose a novel parallelized planning-acting multi-agent framework that significantly enhances the responsiveness and adaptability of LLM-based MAS in dynamic environments. Our framework’s dual-thread architecture with interruptible execution mechanism enables real-time interaction and continuous adaptation, overcoming the limitations of traditional serialized execution paradigms. The comprehensive skill library and automated task decomposition mechanism further improve efficiency and coordination as an engineering contribution. Through extensive experiments, our framework demonstrates superior performance in both collaborative and adversarial scenarios, and achieves more immediate human-agent interaction, highlighting its potential for broader applications.
6 Limitations
While demonstrating significant advantages, our framework has three main limitations. First, the continuous LLM request design leads to higher computational costs, which may limit scalability in resource-constrained environments. Second, LLM hallucinations occasionally result in irrational action interruptions, affecting the reliability of agent behavior. Third, although our framework adapts to multimodal observations, the current fusion method relies solely on visual-language models (VLMs), potentially failing to fully exploit the complementarity of multimodal information and limiting performance in complex scenarios. For future work, we aim to optimize computational efficiency by refining LLM request mechanisms, mitigate hallucination effects through enhanced verification strategies, and improve multimodal fusion techniques to better leverage diverse sensory inputs.
Appendix A Centralized Memory System
The memory of the multi-agent system is stored in JSON files, and agents read and modify their contents by reading and writing to these files. Below is an example of how the memory is stored:
A.1 Agent Observation & Status
"time": "day",
"health": 12.8,
"hunger": 17,
"position": "(10, 65, -8)",
"equipment": [
null,
null,
null,
null,
"bow",
null
"inventory": {
"arrow": 187,
"golden_apple": 4,
"diamond_pickaxe": 1,
"diamond_sword": 1,
"diamond_helmet": 1,
"diamond_chestplate": 1,
"diamond_leggings": 1,
"diamond_boots": 1,
"bow": 1
"nearbyBlocks": [
"end_stone",
"bedrock"
"nearbyEntities": "enderman, ender_dragon",
"nearbyPlayers": [
"Steve",
"Brine",
"Jeb",
"Notch"
A.2 Team Chat Log
"Day 1 - Morning": [
"SystemInfo: Number of End Crystals Remaining: 10",
"SystemInfo: Boss Health: 200",
"SystemInfo: Boss Health: 200",
"Notch: Lets prioritize taking down the end_crystals to prevent the Ender Dragon from healing. I will start attacking the nearest end_crystal with my bow. Everyone, cover me and be ready to assist if needed.",
"Steve: Im ready with my bow and arrows. Lets focus on taking down the end_crystals first to prevent the Ender Dragon from healing. Ill start targeting the nearest end_crystal. Stay alert for any incoming endermen but avoid engaging them unless necessary.",
"SystemInfo: Boss Health: 200",
"Brine: Brine: Lets prioritize destroying the end_crystals to prevent the Ender Dragon from healing. Notch and I will start targeting the nearest crystal. Steve, Alex, and Jeb, please spread out and assist in taking down other crystals. Stay alert for any Endermen but avoid engaging them unless necessary. Report if you need help or spot any crystals.",
"Jeb: Im ready with my bow and arrows. Lets prioritize the end_crystal on the obsidian pillars to prevent the Ender Dragon from healing. Ill start targeting the nearest crystal. Stay alert for any incoming endermen but avoid engaging unless necessary."
A.3 Team Observation Log
"obs_entries": [
"Notch: Team, Im fully equipped with diamond armor, a bow, and 200 arrows. I see an end_crystal and an ender_dragon nearby, along with some endermen. Brine, Jeb, Steve, and Alex are also close. Lets focus on taking down the end_crystal first to prevent the Ender Dragon from healing. Everyone, prepare your bows and stay alert for any incoming endermen. Ready when you are!",
"Brine: Team, Im currently well-equipped with a bow, 200 arrows, and diamond armor. The ender_dragon is nearby but hasnt attacked yet. I see Steve, Notch, Alex, and Jeb around. Lets focus on taking out the end_crystals first. I suggest we spread out to cover more ground and target the crystals from a distance. Stay alert for any Endermen, but avoid engaging them unless necessary. Report if you spot any crystals or need assistance.",
"Jeb: All end_crystals have been destroyed. Focusing on the Ender Dragon now. Everyone, keep your distance and use bows to attack from a safe range. Stay alert for any incoming endermen but prioritize the Ender Dragon. Lets finish this!",
"Steve: Im in critical condition and need immediate assistance! My health is down to 0. Brine, Notch, Jeb, please cover me or help me get to safety. I have a full bow and golden apples but cant fight effectively in this state. Endermen are nearby, be cautious!",
A.4 Team Action Log
"Notch": [
[…]
"Steve": ...
Part of above information is provided as human prompts for agents to plan actions. In addition to this, agents are also provided with identity information such as their name, team affiliation, and teammates’ names, to simulate a human-like multi-agent collaborative scenario.
Appendix B Minecraft Comprehensive Skill Library
Our proposed skill library is implemented based on MineflayerPrismarineJS (2023).To simulate the process of solving prerequisite steps in the actual game to eventually obtain specific items, this skill library has introduced a recursive task decomposition mechanism. From the implementation level, all interfaces can recursively invoke each other or themselves and pass parameters.
B.1 Basic Skill Interfaces
  • obtainItem(bot, count, type): Automates the collection of items.
  • mineItem(bot, count, type, explore_direction, explore_time): Mines using specific tools. If the exploration parameters are not specified, the exploration direction is chosen randomly. If the item type is a subterranean mineral like diamond, the exploration direction is set to down (0, -1, 0). The exploration time limit defaults to five minutes (6000 ticks in Minecraft).
  • craftItem(bot, count, type, need_crafting_table): Crafts items. If no parameter is specified, the crafting is done using a crafting table by default.
  • smeltItem(bot, count, type, fuel): Smelts or cooks items in a furnace. If no parameters are provided, coal is used as the default fuel.
  • collectItem(bot, count, type, function): Collects items by killing animals or using other special methods.
  • chatMessage(bot, message, team_name): Sends a chat message to the team.
  • getItemFromChest(bot, chest_position, items_to_get): Retrieves specific items from a chest at a given location.
  • depositItemIntoChest(bot, chest_position, items_to_deposit): Deposits specific items into a chest at a given location.
  • combatWithEntity(bot, mob_name, weapon, loop): Automatically equips the highest quality armor and weapon from the inventory to fight with an entity (e.g., animals or hostile mobs). If no parameters are specified, the default weapon is a sword. The loop parameter determines whether to continue fighting the same type of entity.
  • combatWithPlayer(bot, player_name, weapon): Automatically equips the highest quality armor and weapon from the inventory to fight with a specific player. If no parameters are provided, the default weapon is a sword.
  • initialInventory(bot, item_dict): Initializes the player’s inventory with specified items.
  • equipBestToolOrArmor(bot, type): Automatically equips the highest quality tool, weapon, or armor of the specified type.
  • listenChat(bot, player_name): Continuously listens to a player’s chat messages.
B.2 Recursive-related Interfaces
  • preTool(item): Retrieves the minimum prerequisite tools required to gather a specific item.
  • preItem(item): Retrieves the prerequisite items required to craft a specific item and indicates whether a crafting table is necessary.
  • preSmelt(item): Retrieves the prerequisite items required to smelt a specific item.
  • preCollect(item): Retrieves the prerequisite entities required to collect a specific item.
  • getFunc(item): Retrieves the method used to collect a specific item.
This library is scalable and can be easily extended to accommodate new items, recipes, and tasks. Currently, it supports the collection of over 790 types of items across all three dimensions of Minecraft (Overworld, Nether, and End). However, the library does not yet support some specialized gameplay features. Future updates will aim to expand the library’s functionality, adding support for more complex tasks.
Appendix C Experiment Details
Experiments were conducted using the Qwen-Plus model API provided by Alibaba Cloud. In the multiplayer local area network (LAN) server of Minecraft, agents connected to the game via different ports and interacted with the environment and other players.
C.1 Resource Collection Task
We implemented a comprehensive Minecraft resource collection benchmark, which allows multi-agent collaboration to automatically collect resources based on a given requirement dictionary, making this benchmark general and scalable. We conducted experiments on the following representative resource collection tasks:
Task Definitions:
  • Iron Tool Set: { ’iron_pickaxe’: 1, ’iron_shovel’: 1, ’iron_hoe’: 1, ’iron_axe’: 1 } — A set of commonly used iron tools in Minecraft.
  • Diamond Armor: { ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1 } — A full set of diamond armor in Minecraft.
  • Redstone Devices: { ’repeater’: 1, ’piston’: 1, ’dropper’: 1 } — Common redstone components and devices in Minecraft.
  • Navigation Kit: { ’compass’: 1, ’clock’: 1, ’map’: 1 } — A set of tools commonly used for navigation in Minecraft.
  • Transport System: { ’minecart’: 1, ’rail’: 16, ’powered_rail’: 6 } — Minecarts and tracks used to complete a transportation system.
  • Food Supplies: { ’beef’: 1, ’chicken’: 1, ’porkchop’: 1 } — Common animal meat food items in Minecraft.
  • Building Materials: { ’stone_bricks’: 4, ’glass’: 4, ’iron_door’: 1 } — Building materials commonly used for constructing doors, windows, and walls.
  • Storage System: { ’hopper’: 1, ’chest’: 1, ’barrel’: 1 } — Tools commonly used for storing items in Minecraft.
Each round of the experiment is conducted in a world generated with a random seed to ensure the generalizability of the results.
C.2 Boss Combat Task
We manually implemented three representative and challenging task scenarios: defeating the Elder Guardian in an ocean monument in the Overworld, defeating the Wither in the Nether, and defeating the Ender Dragon in the End. Defeating the Ender Dragon represents the final achievement in the game. Prior to each combat, agents were equipped with standardized combat resources, including a full set of diamond armor (validated as efficiently collectible through our resource collection experiments), bow and arrows, and some consumables. In all experiments, the standardized combat resources remained consistent. A single LLM call initialized resource allocation, distributing equipment to each agent through a dictionary-based assignment system. Agents were then teleported to designated combat locations, where boss monsters and supporting entities (e.g., Guardians near the Elder Guardian, End Crystals for Ender Dragon healing) were spawned to ensure scenario complexity. During combat, real-time progress information (boss health, remaining End Crystals, etc.) was communicated through the chat system to support strategic planning. Task completion time was measured excluding initialization phases.
Evaluation metrics are defined as follows: health ratio = $\frac{1}{n}\sum_{i=1}^{n}\frac{h_{i}}{H_{\text{max}}}\times 100\%$ ($h_{i}$: agent’s remaining health, $H_{\text{max}}$: agent’s max health) and progress = $\frac{H_{\text{boss}}-h_{\text{boss}}}{H_{\text{boss}}}\times 100\%$ ($h_{\text{boss}}$: boss’s remaining health, $H_{\text{boss}}$: boss’s max health).
C.3 Adversarial PVP Task
Similar to the Boss Combat Task, in this task, agents were provided with combat resources. Results were recorded when one team of agents was entirely defeated. After multiple experiments, victory rates and other relevant metrics were calculated. In order to promote fair confrontation, we give each team exactly the same initial supplies. Here is an example of initial inventory given to an agent : { ’wooden_sword’: 1, ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1, ’bow’: 1, ’arrow’: 200, ’golden_apple’: 20 }
C.4 Human-Agent Interaction
This section presents examples of human-agent interactions in dynamic environments, highlighting how agents and human players collaborate in various tasks.
Scenario 1: Human-Agent Collaboration in Resource Collection Tasks
Agent Steve’s Environment Information:
"time": "day",
"health": 20,
"hunger": 20,
"position": "(-223, 60, -159)",
"equipment": [
null,
null,
null,
null,
"crafting_table",
null
"inventory": {
"crafting_table": 1
"nearbyBlocks": [
"dirt",
"grass_block",
"oak_log",
"oak_leaves"
"nearbyEntities": "pig",
"nearbyPlayers": ["Alex", "Notch"]
Human Player’s Chat Message:
’Hey Steve, I’ve got some beef, can you hunt a pig for obtaining porkchops?’
Agent Steve’s Planning Action:
"skill": "obtainItem(bot, 10, \’porkchop\’)",
"interrupt": false,
"reason": "Following the human players request to hunt a pig for obtaining porkchops. Since there are pigs nearby, I will proceed with hunting to gather porkchops."
Scenario 2: Human Player Directing Agents for Combat Tasks
Agent Steve’s Environment Information:
"time": "day",
"health": 1.2,
"hunger": 17,
"position": "(6, 66, 13)",
"equipment": [
"diamond_helmet",
"diamond_chestplate",
"diamond_leggings",
"diamond_boots",
"bow",
null
"inventory": {
"arrow": 189,
"golden_apple": 4,
"wooden_sword": 1,
"bow": 1
"nearbyBlocks": [
"end_stone",
"bedrock"
"nearbyEntities": "ender_dragon",
"nearbyPlayers": [
"Steve",
"Notch",
"Jeb"
Human Player’s Chat Message:
’Steve, you are in grave danger. Cease your attack immediately and prepare to restore health!’
Agent Steve’s Planning Action:
"skill": "consumeItem(bot, \’golden_apple\’, false)",
"interrupt": true,
"reason": "My health is critically low at 1.2, and I need to restore it immediately to continue fighting the ender_dragon effectively. I will follow the human players advice and prioritize my health."

### v2

@@
-Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems
-Abstract
+Parallelized Planning-Acting for Multi-Agent LLM Systems in Minecraft
+Abstract.
 Recent advancements in Large Language Model (LLM)-based Multi-Agent Systems (MAS) have demonstrated remarkable potential for tackling complex decision-making tasks.
 However, existing frameworks inevitably rely on serialized execution paradigms, where agents must complete sequential LLM planning before taking action.
-This fundamental constraint severely limits real-time responsiveness and adaptation, which is crucial in dynamic environments with ever-changing scenarios.
+This fundamental constraint severely limits real-time responsiveness and adaptation, which is crucial in dynamic environments with ever-changing scenarios like Minecraft.
 In this paper, we propose a novel parallelized planning-acting framework for LLM-based MAS, featuring a dual-thread architecture with interruptible execution to enable concurrent planning and acting.
 Specifically, our framework comprises two core threads: (1) a planning thread driven by a centralized memory system, maintaining synchronization of environmental states and agent communication to support dynamic decision-making; and (2) an acting thread equipped with a comprehensive skill library, enabling automated task execution through recursive decomposition.
-Extensive experiments on challenging Minecraft demonstrate the effectiveness of the proposed framework.
-Parallelized Planning-Acting for Efficient LLM-based Multi-Agent Systems
-[…]
-[…]
-1 Introduction
-Figure 1: An overview of the proposed multi-agent system
+Extensive experiments on Minecraft demonstrate the effectiveness of the proposed framework.
+Key words and phrases:
+Multi-Agent Systems; Large Language Models
+doi:
+EXAJ9853
+11footnotetext: Corresponding author.
+1. Introduction
 Multi-Agent Systems (MAS) have become a well-established paradigm for tackling complex decision-making problems Hong et al.
 (2023); Chen et al.
-(2024b); Dong et al.
+(2024c); Dong et al.
 (2024), with early efforts primarily relying on reinforcement learning Busoniu et al.
@@
 (2023); Qwen et al.
+(2024); Yang et al.
 (2025); Dubey et al.
-(2024) has since revolutionized MAS by adding natural language understanding and generation capabilities, enabling agents to engage in more sophisticated collaboration.
+(2024); Kimi (2025) has since revolutionized MAS by adding natural language understanding and generation capabilities, enabling agents to engage in more sophisticated collaboration.
 LLMs have significantly enhanced the flexibility and versatility of MAS, opening the door to more complex tasks and dynamic interactions in real-world applications Wu et al.
@@
 (2024); Chen et al.
-(2024c); AL et al.
+(2024a); AL et al.
 (2024).
 Recent works have demonstrated the potential of LLM-based MAS in various domains.
-MetaGPT Hong et al.
-(2023) introduces an innovative meta-programming framework to enhance task decomposition and agent collaboration, AgentVerse Chen et al.
-(2024b) improves collaborative performance by orchestrating expert agents, and VillagerAgent Dong et al.
+AgentVerse Chen et al.
+(2024c) improves collaborative performance by orchestrating expert agents, and VillagerAgent Dong et al.
 (2024) tackles task dependencies in complex environments using DAG-based task decomposition.
-Despite these advancements, most current frameworks still rely on serialized execution, where planning and acting occur sequentially for each agent.
-This serialized nature creates a substantial bottleneck when handling dynamic information, particularly evident in dynamic settings like Minecraft, where the environment is constantly changing, even without the intervention of agents.
-Although Voyager Wang et al.
-(2023) pioneered LLMs-based agents in Minecraft, its approach, which pauses the game server during LLM interactions, effectively staticizes the dynamic environment, thus failing to fully address the real-time responsiveness needed in such settings.
-Subsequent works in Minecraft agents Liu et al.
-(2024); Zhao et al.
-(2024) have followed this approach, but they also inherit the same limitations.
+Despite these advancements, most current frameworks applied in dynamic environments still rely on serialized execution, where planning and acting occur sequentially for each agent.
+This serialized nature creates a substantial bottleneck when handling dynamic information, particularly evident in dynamic settings like Minecraft, a game that features a vast and diverse world with various terrains, resources and creatures.
+Such an environment serves as an ideal testbed for evaluating the capabilities of MAS in open-world scenarios due to its constant environmental changes and rich interaction possibilities.
+While Voyager Wang et al.
+(2023) pioneered LLM-based agents in Minecraft, it relies on pausing the game server during planning to staticize the environment for the agent in order to prevent interference from environmental changes.
+However, in real-world dynamic environments, especially in multi-agent systems, it is not feasible to halt the actions of other agents while one agent is planning.
+This limitation hinders real-time interaction and reduces the system’s adaptability to sudden environmental changes or incoming information from other agents during LLM invocations.
 Our analysis reveals three critical challenges in current LLM-based MAS for dynamic environments.
 First, inflexible action scheduling is prevalent, as many existing agent frameworks rely on serialized execution, requiring agents to wait for a language model response before proceeding with further actions.
-This rigidity complicates the handling of unexpected environmental changes; Second, limited replanning capabilities hinder agents’ performance, as they often execute actions to completion without interruption.
+This rigidity complicates the handling of unexpected environmental changes.
+Second, limited replanning capabilities hinder agents’ performance, as they often execute actions to completion without interruption.
 This lack of adaptability prevents agents from effectively reconsidering or adjusting their plans in response to urgent and unforeseen events, diminishing their overall effectiveness.
@@
 This results in delayed observational data sharing, causing agents to operate based on outdated information, which in turn limits the team’s coordination and efficiency.
-In this paper, we propose a parallelized planning-acting framework that introduces a dual-thread architecture with interruptible execution for efficient LLM-based MAS in dynamic environments, as shown in Fig.
-1.
+In this paper, we propose a parallelized planning-acting framework that introduces a dual-thread architecture with interruptible execution for efficient LLM-based MAS in dynamic environments.
 Our architecture decouples LLM reasoning from action execution, enabling concurrent planning and acting.
@@
 Our core contributions are summarized as follows:
-  • We propose a parallelized planning-acting framework that decouples planning and acting into parallelized dual threads with interruptible execution for efficient LLM-based MAS.
-  • We develop a centralized memory system to support the planning thread, ensuring that agent decisions are always informed by the latest environmental changes and interactions.
+  • We propose a parallelized planning-acting framework that decouples planning and acting into dual threads with interruptible execution for efficient LLM-based MAS.
+  • We develop a centralized memory system to support the planning thread, ensuring agent decisions are informed by the latest environmental changes and interactions.
   • We design a comprehensive skill library to empower the acting thread, enabling efficient task execution through recursive task decomposition.
-  • Experimental results on Minecraft demonstrate a paradigm shift from serialized deliberation to parallelized interaction, yielding notable improvements in efficiency and coordination.
-2 LLM-based Multi-Agent Framework
-We propose a novel parallelized planning-acting framework based on LLMs, designed to enable collaboration and decision-making among agents in complex dynamic environments.
-Our framework introduces three key innovations: (1) A dual-thread architecture with an interruptible execution mechanism, enabling concurrent planning and acting, (2) A real-time updated centralized memory system supporting the planning thread, ensuring that agents’ decisions are informed by the latest environmental changes and team interactions, and (3) A Minecraft comprehensive skill library supporting the acting thread, automating task execution by proposing a recursive task decomposition mechanism.
-While the skill library represents an engineering contribution, it is important to note that our framework is generic and can be extended to other domains with minimal adaptation.
-This flexibility highlights the broader applicability of our approach, while Minecraft provides a rich and accessible testbed for evaluating its performance in dynamic environments.
-2.1 Parallelized Planning-Acting Framework
-Inspired by the human ability to think and act simultaneously, our framework features a dual-thread architecture in Fig.
-2 that separates planning (supported by LLMs and the centralized memory system) and acting (executed by the comprehensive skill library).
-Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{n}\}$ denote the set of agents.
-Our framework operates as follows:
+  • Experimental results on Minecraft demonstrate a paradigm shift from serialized deliberation to parallelized interaction, yielding notable improvements in efficiency, coordination and adaptability in dynamic environments.
+2. LLM-based Multi-Agent Framework
+We propose a novel parallelized planning-acting framework for LLM-based multi-agent systems as shown in Fig.
+1, designed to support real-time inter-agent collaboration in dynamic scenarios such as the open-world environment of Minecraft.
+Our framework introduces three key innovations: (1) A dual-thread architecture with an interruptible execution mechanism, enabling concurrent planning and acting, (2) A real-time updated centralized memory system supporting the planning thread, ensuring that agents’ decisions are informed by the latest environmental changes and team communications, and (3) A comprehensive skill library supporting the acting thread, automating task execution by proposing a recursive task decomposition mechanism.
+Figure 1.
+An overview of our parallelized planning-acting framework.
+In the multi-agent system, each agent operates with independent planning and acting threads in parallel.
+Planning threads are supported by a centralized memory system for decision-making, while acting threads utilize a DAG-based comprehensive skill library for task execution.
+After each planning step, the latest planned action is stored in the action buffer, and the agent performs a communication.
+Whenever an action is completed or the interrupt mechanism is triggered, the agent will continue to execute the skill function in the action buffer.
+2.1. Parallelized Planning-Acting Framework
+Inspired by the human ability to think and act simultaneously, our framework adopts a dual-thread architecture (Fig.
+1) that decouples planning (driven by LLMs and the centralized memory system) from acting (executed by a comprehensive skill library).
+Let $\mathcal{G}=\{g_{1},g_{2},\dots,g_{n}\}$ denote the set of agents.
+The planning and acting threads operate asynchronously and independently, communicating only through a shared action buffer.
+  • Planning Thread: The planning thread continuously monitors the environment and generates new action proposals for agent $g_{i}$ based on the system prompt $S$, the agent’s current observation $O_{i}$, the latest team chat logs $C$, and its current action $A_{i}$.
+At any time, the LLM may propose a new action together with an interruption flag:
+    $A_{i}^{\text{new}},flag_{intr}=\text{LLM}(S,O_{i},C,A_{i}).$ (1)
+    This proposed action $A_{i}^{\text{new}}$ is then written into a shared action buffer, which acts as a communication channel between the planning and acting threads.
+The buffer is implemented as a single-slot queue: if it is already occupied, the previous action will be overwritten.
+This ensures that the buffer always holds the most up-to-date action recommendation from the planner, reflecting the latest context.
+The interruption mechanism is also completely controlled by the LLM: it may trigger a restart if the new action is judged more urgent, or if the current action is no longer meaningful.
+When $flag_{intr}=\text{True}$, the planner issues a restart signal to the acting thread, ensuring that ongoing execution will be terminated and replaced by the new plan.
+  • Acting Thread: The acting thread is responsible for executing skills from the comprehensive skill library.
+Let $A_{i}^{\text{curr}}$ denote the action currently being executed, and $A_{i}^{\text{new}}$ the action in the buffer.
+Upon restart, the acting thread immediately aborts its ongoing skill execution and fetches the latest action from the buffer.
+Otherwise, it continues executing $A_{i}^{\text{curr}}$ until completion, while periodically checking for updates.
+This design makes the acting thread reactive to planner-issued interrupts, rather than making its own decisions about preemption.
+This decoupled design allows flexible, interruptible execution: planners can revise intentions at high frequency, while actors respond dynamically by replacing or preempting ongoing actions.
+Compared to fixed scheduling, this architecture significantly enhances responsiveness and adaptability in dynamic environments.
+Algorithm 1 summarizes the procedure.
+Algorithm 1 Parallelized Planning-Acting Framework
+1:Initialize agent set $\mathcal{G}\leftarrow\{g_{1},\dots,g_{n}\}$
 […]
-    $A_{i}^{t+1}=\text{LLM}(S,O_{i}^{t},C^{t},A_{i}^{t}).$ (1)
-    The planned action $A_{i}^{t+1}$ is then written into an action buffer, which serves as the communication channel between the planning thread and the acting thread.
-This action buffer is implemented as a single-slot queue, allowing the planning thread to write and the acting thread to read.
-If the buffer is already occupied, the previous action is discarded to make space for the new one.
-This mechanism ensures that the planning thread always places the most up-to-date action (based on the latest observations and agent status) into the buffer, while the acting thread always retrieves the most current action, never operating outdated plans.
-  • Acting Thread: The acting thread retrieves actions from the action buffer and executes them according to the following rule:
-    $s_{i}=\begin{cases}A_{i}^{t+1}&\text{if }p(A_{i}^{t+1})>p(A_{i}^{t}),\\ A_{i}^{t}&\text{otherwise}.\end{cases}$ (2)
-    Here, $s_{i}$ represents the skill execution of agent $a_{i}$, $A_{i}^{t}$ is the current action being executed, $A_{i}^{t+1}$ is the new action in the buffer, and $p$ is the priority of an action.
-If an interruption is triggered—i.e., when the LLM determines that the new action has a higher priority than the current action—the planning thread sends an interrupt signal to the acting thread.
-The acting thread immediately exits and restarts with the new action.
-If no interruption is triggered, the acting thread completes the current action $A_{i}^{t}$ before retrieving and executing the next action $A_{i}^{t+1}$ from the buffer.
+[…]
+4:procedure PlanningThread($g_{i}$)
+5:  while not $\text{TaskCompleted}()$ do
+6:   O_i $\leftarrow$ GetObservation(g_i)
+7:   C $\leftarrow$ M.GetChatLogs()
+8:   (new_action, intr) $\leftarrow$ LLM(S, O_i, C, cur_action)
+9:   if $\texttt{intr}=True$ then
+10:     Restart(ActingThread($g_{i}$))
+11:   end if
+12:   $Q.\text{put}(\texttt{new\_action},\text{overwrite=True})$
+13:  end while
+14:end procedure
+15:procedure ActingThread($g_{i}$)
+16:  while not $\text{TaskCompleted}()$ do
+17:   if $Q.\text{has\_new}()$ then
+18:     new_action $\leftarrow Q.\text{get}()$
+19:     cur_action $\leftarrow$ new_action
+20:   end if
+21:   if $\texttt{cur\_action}\neq\text{None}$ then
+22:     ExecuteSkill(cur_action)
+23:     $M.\text{UpdateExecution}(\texttt{cur\_action})$
+24:   end if
+25:  end while
+26:end procedure
+27:Start PlanningThread and ActingThread for each $g_{i}\in\mathcal{G}$
+28:Wait until all threads terminate
 Latency Analysis.
@@
 The latency reduction $\Delta T$ can be expressed as:
-$\Delta T\approx\sum_{k=1}^{n}\left(T_{\text{plan}}^{(k)}+T_{\text{act}}^{(k)}-\max(T_{\text{plan}}^{(k)},T_{\text{act}}^{(k)})\right).$ (5)
-This analysis highlights two key advantages of our framework: 1) The initial planning latency $T_{\text{plan}}^{(1)}$ is effectively amortized over subsequent actions, and 2) The overlapping of planning and acting phases successfully conceals $T_{\text{plan}}$ when $T_{\text{act}}>T_{\text{plan}}$ (our proposed comprehensive skill library in Section 2.3 ensures this condition is well-maintained).
-Our experimental results in Section 3.4 demonstrate the overall efficiency advantage of the parallelized framework.
-By decoupling planning and acting into parallelized threads, our framework strikes a balance between efficiency and flexibility, allowing agents to respond dynamically to unpredictable environments with enhanced adaptability and performance.
-2.2 Centralized Memory System
-Figure 2: An overview of the proposed parallelized planning-acting architecture
+$\begin{split}\Delta T&\approx\sum_{k}^{n}\left(T_{\text{plan}}^{(k)}+T_{\text{act}}^{(k)}-\max\left(T_{\text{plan}}^{(k)},T_{\text{act}}^{(k)}\right)\right)\\ &=\sum_{k}^{n}T_{\text{plan}}^{(k)}\quad\text{if for all }k,\,T_{\text{act}}^{(k)}>T_{\text{plan}}^{(k)}\end{split}$ (4)
+This analysis highlights that the overlapping of planning and acting phases successfully conceals $T_{\text{plan}}$ when $T_{\text{act}}>T_{\text{plan}}$.
+We propose the comprehensive skill library in Section 2.3 to ensure this condition is well-maintained.
+For instance, a complex long-duration skill that might take several minutes to execute can be recursively decomposed and automated, while LLM reasoning typically only requires a few seconds.
+It is important to emphasize that our primary goal is to enable real-time and dynamic interaction, while latency reduction is a secondary benefit.
+Fig.
+2 vividly presents a compact timeline diagram that illustrates the overlap between planning and acting, along with a concrete interrupt example.
+$2.$ Centralized Memory System
 To facilitate effective coordination, we implement a centralized memory system $M$ that stores and manages information at the team level. The memory is updated at each time step $t$ as follows:
@@
 This unified repository enables agents to access and utilize relevant information during task execution, ensuring efficient team coordination:
-  • Observation Records: Each agent’s observations are continuously updated in the centralized memory, reflecting the latest agent status and environmental state.
+Figure 2.
+Schematic diagram of the timeline for parallel operation of the planning-acting dual threads.
+Among them, plans 2-3 are deemed not urgent enough to trigger the interrupt mechanism and thus are not actually executed.
+After the action 1 corresponding to plan 1 is completed, the latest action 4 corresponding to plan 4 is executed directly.
+Plan 6 sends an interrupt signal to suspend action 4, and the execution of action 6 starts immediately.
+  • Observation Records: Each agent’s observations are continuously updated in a polling manner (e.g.
+update per second), reflecting the latest agent status and environmental state.
 These observations are associated with the respective agent, allowing the team to maintain a comprehensive and up-to-date view of the environment.
-  • Chat Logs: All team chat messages are stored in the centralized memory, with long-term retention to support historical analysis and decision making.
-During planning, agents can retrieve the most recent chat messages to incorporate team insights into their strategies.
-This ensures that decisions are informed by the collective knowledge of the team.
-  • Action History: The centralized memory also records the actions taken by each agent, providing a detailed history of task execution.
-These logs are useful for behaviours refining and performance analysis.
-We implement two types of multi-agent communication (passive and active) ensuring that the centralized memory remains a dynamic resource for team coordination:
-  • Passive Communication: In the planning thread, after each planning cycle, the LLM generates a chat message based on the agent’s latest observations, which is then sent to the centralized memory’s chat logs.
-This ensures that passive communication, reflecting updated observations, can run concurrently with action execution.
+  • Chat Logs: Team chat messages are always updated in real time, with long-term retention to support historical analysis and decision making.
+During planning, agents can retrieve the most recent messages to incorporate team insights into their strategies, ensuring team collaboration.
+  • Action History: Actions taken by each agent are also recorded, providing a detailed history of task execution.
+During planning, agents need to make decisions based on their current action and determine whether to interrupt it.
+We implement two types of multi-agent communication, including passive communication and active communication, ensuring a dynamic and diverse resource for team coordination among agents:
+  • Passive Communication: In the planning thread, the LLM generates a chat message based on the agent’s latest observations after planning, which is then sent to the centralized memory’s chat logs.
+This ensures that updated observations can run concurrently with action execution.
 While an agent is performing actions, its observations are continuously updated and shared with the team, enabling real-time coordination based on the most current environmental information.
-  • Active Communication: In the acting thread, agents can actively choose to send chat messages by performing a chat action.
-This action allows the agent to share any information with teammates, updating the chat logs in real-time.
-This form of communication ensures that agents can respond dynamically and share critical information during action execution, facilitating efficient and up-to-date information exchange between agents.
-2.3 Comprehensive Skill Library
-To enable seamless interaction between agents and the Minecraft environment, we develop a comprehensive skill library that encapsulates a wide range of in-game actions.
+  • Active Communication: In the acting thread, agents can actively choose to send chat messages by performing a chat action implemented by the comprehensive skill library.
+This allows the agent to share any information with teammates, updating the chat logs in real-time.
+This form of communication ensures that agents can respond dynamically and share critical information, rather than passively communicating after the current action is over.
+By supporting both passive and active communication modes, this system effectively addresses the challenge of memory sharing latency, ensuring that agents always operate on the latest team knowledge.
+For concrete examples illustrating these concepts, please refer to Appendix.
+2.3. Comprehensive Skill Library
+To enable seamless interaction between agents and the Minecraft environment, we develop a comprehensive skill library based on Mineflayer PrismarineJS (2023) that encapsulates a wide range of in-game actions.
 The library provides high-level APIs for tasks such as resource collection, combat, exploration, and communication.
-For further technical details, please refer to Appendix B.
-A key feature of the comprehensive skill library is the implementation of a recursive task decomposition mechanism, which automates the completion of prerequisite tasks such as mining raw materials and crafting necessary tools.
+For further technical details, please refer to Appendix.
+Our comprehensive skill library implements a recursive task decomposition mechanism, which automates the completion of prerequisite tasks such as mining raw materials and crafting necessary tools.
 This automation ensures that agents can perform complex resource collection tasks with minimal manual intervention, enabling the automated collection of over 790 types of items in Minecraft, surpassing all existing methods Wang et al.
@@
     where $t_{i}\in\mathcal{T}$ is the target item type (All collectible items in Minecraft), $c_{i}\in\mathbb{N}^{+}$ denotes required quantity, and $f_{i}$ specifies the operation type for obtaining the item.
-  • Edge set $E\subseteq V\times V$ encodes prerequisite task dependencies:
+  • Edge set $E\subseteq V\times V$ encodes task dependencies:
     $(v_{j},v_{i})\in E\iff v_{j}\in\text{pre}(v_{i}),$ (7)
-    where $\text{pre}(v_{i})$ gives the prerequisite tasks for producing $t_{i}$.
-  • Weight function $\phi:E\to\mathbb{Q}^{+}$ defines material conversion rates:
+    where $\text{pre}(v_{i})$ gives the prerequisite tasks of $t_{i}$.
+  • Weight function $\phi:E\to\mathbb{Q}^{+}$ defines conversion rates:
     $\phi(v_{j},v_{i})=\frac{r_{ij}}{n_{\text{out}}},$ (8)
-    with $r_{ij}$ being the required quantity of $t_{j}$ per operation, and $n_{\text{out}}$ being the output quantity per operation.
+    with $r_{ij}$ being the required quantity of $t_{j}$ and $n_{\text{out}}$ being the output quantity per operation.
 The recursive resolution process follows:
 $\Psi(v_{i})=\bigcup_{(v_{j},v_{i})\in E}\left\{\Psi\left(v_{j}^{\left(\phi(v_{j},v_{i})\cdot c_{i}\right)}\right)\right\}\cup\{v_{i}\},$ (9)
-where $v_{j}^{(k)}$ denotes a task requiring $k$ units of $t_{j}$, and the base case $\Psi(v_{i})=\emptyset$ applies when $I(t_{i})\geq c_{i}$, with $I$ representing the current inventory state, which means the task is accomplished.
-3 Experiments
-3.1 Experiment Setup
-We aim to address the following questions through further experimentation: (1) Can our comprehensive skill library enable agents to efficiently accomplish basic tasks within Minecraft’s open-world environment?
-(2) How well can our proposed multi-agent framework perform on highly challenging and complex tasks?
-(3) Can our proposed parallelized planning-acting framework better support multi-agent collaboration, competition, and human-agent interaction?
-All experiments are conducted in the Minecraft Java Edition version 1.19.4 gaming environment using the Qwen-Plus model  Qwen et al.
+where $v_{j}^{(k)}$ denotes a task requiring $k$ units of $t_{j}$, and the base case $\Psi(v_{i})=\emptyset$ applies when $I(t_{i})\geq c_{i}$, with II representing the current inventory state, which means the task is accomplished.
+This recursive task decomposition mechanism effectively models task prerequisite relationships as a Directed Acyclic Graph (DAG), where complex tasks are dynamically decomposed into atomic subtasks through automated dependency resolution.
+Unlike conventional approaches that require explicit step-by-step navigation from initial states to target objectives, our implementation enables agents to directly invoke high-level skills while the system automatically handles the recursive resolution of all prerequisite conditions.
+This design allows our multi-agent system to remain efficient by offloading detailed task execution to the skill library while leveraging LLMs for strategic decision-making and dynamic prioritization.
+Moreover, the recursive task decomposition mechanism can be generalized to other environments beyond Minecraft.
+It is intuitive that allowing an LLM to independently plan the task dependencies embedded within a DAG would not be as efficient as utilizing a fixed mechanism.
+Consider a Directed Acyclic Graph (DAG) $\mathcal{G}=(V,E)$ with $n=|V|$ nodes and $e=|E|$ edges, representing task dependencies.
+Let the shortest path length from the source node $v_{s}$ (a node with zero in-degree) to a target node $v_{t}$ (a node with zero out-degree) be denoted as $L(v_{s},v_{t})$.
+This path corresponds to the minimal sequence of prerequisite tasks that must be executed to achieve the final objective represented by $v_{t}$.
+In traditional planning approaches where each task is sequentially inferred by an LLM, the number of required model calls is at least $L(v_{s},v_{t})+1$, i.e., one for each node along the critical path.
+In contrast, our mechanism resolves all prerequisite dependencies automatically through the DAG structure, requiring only a single LLM invocation for the high-level task $v_{t}$.
+This significant reduction in LLM usage not only improves computational efficiency but also reduces potential error propagation across multiple reasoning steps.
+Therefore, our methodology offers a generalizable framework for integrating LLMs with structured task execution systems across diverse domains beyond Minecraft.
+The comprehensive skill library also demonstrates strong scalability and can be readily extended to accommodate new updates.
+Newly introduced skills can be seamlessly integrated by simply updating the prerequisite dependency DAG that encodes task relationships without new interfaces development or extensive code modifications.
+This modular design ensures continued functionality with minimal maintenance effort, making the library both future-proof and adaptable to evolving domain requirements.
+3. Experiments
+3.1. Benchmark Task Design
+Our experiments are designed to validate the framework’s capabilities, leveraging Minecraft as a testbed while focusing on general Multi-Agent System (MAS) competencies.
+Existing Minecraft agent methods suffer from a lack of a comprehensive skill library, which hinders their ability to learn complex strategies for tackling challenging tasks.
+Consequently, we have developed more demanding tasks built upon the existing evaluation paradigm, rather than limiting our scope to basic tasks solely for baseline comparisons.
+For example, defeating the Ender Dragon is widely recognized as Minecraft’s ultimate challenge, which remains unachievable with current methods to the best of our knowledge, serves as one of our evaluation tasks.
+Our benchmark comprises a diverse range of tasks and supports standardized evaluation for potential future research.
+Refer to the Appendix for more details.
+  • Resource Collection Task: Evaluates the foundation of our framework through fundamental mining and crafting tasks in Minecraft.
+These tasks aim to validate the effectiveness of our comprehensive skill library by requiring agents to complete compound items with deep dependency chains and measure multi-agent coordination efficiency in distributed resource collection.
+we define a set of representative resource collection tasks, such as Diamond Armor serves as a foundation for challenging combat scenarios in the game.
+  • Boss Combat Task: Assesses dynamic adaptation and strategical coordination through combat scenarios against powerful bosses.
+In Minecraft, there are three primary world dimensions: Overworld, Nether, and End.
+Each dimension hosts powerful boss monsters whose defeat is considered pinnacle challenges for players of Minecraft.
+Based on this, we define three representative combat tasks.
+  • Adversarial PVP Task: Compares two frameworks directly by having them engage in direct combat scenarios. In this task, two teams of agents engage in battles with each other.
+3.2. Experimental Setup
+All experiments are conducted in the Minecraft gaming environment using the Qwen-Plus model Qwen et al.
 (2024), while multi-modal experiments utilized the Qwen-VL-Plus model Bai et al.
 (2023).
-The game server operates continuously without pausing during interactions with LLMs, thereby necessitating that all agents perform in real-time.
-All evaluation metrics are captured dynamically during these interactions.
-Metrics such as ’time’ denote the total duration required by the multi-agent systems to complete their designated tasks, encompassing the time expended on communications with LLMs.
-3.2 Benchmark Task Scenarios
-To evaluate the generalizability of our multi-agent framework in Minecraft, we designed three distinct task scenarios: Resource Collection, Boss Combat, and Adversrial Player versus Player (PvP).
-These scenarios were selected to cover a range of application domains, showcasing the versatility of our multi-agent framework across diverse contexts.
-  • Resource Collection: This task involves gathering resources in the Minecraft world, evaluating the efficiency of the comprehensive skill library and the coordination of multi-agent system in performing repetitive and structured tasks.
-  • Boss Combat: This task requires the multi-agent system to defeat a strong boss-level enemy in a predefined, highly intricate combat environment, evaluating advanced strategic planning, real-time decision-making, effective collaboration among agents and the system’s ability to handle dynamic and challenging environments.
-  • Adversarial PVP: This task focuses on team-based adversarial competition, where two teams of agents compete against each other in a combat scenario.
-The primary objective is to evaluate the agents’ ability to strategize, adapt, and outperform opponents.
-3.3 Main Results
-3.3.1 Resource Collection Task
-We first evaluate the effectiveness of our skill library’s recursive task decomposition mechanism by assessing a single agent’s performance on basic resource collection tasks in Minecraft.
-This initial evaluation aimed to verify the comprehensive skill library’s capability in automating complex workflows, thereby validating its efficiency and reliability in handling fundamental Minecraft tasks.
-As shown in Table 1, when the recursive task decomposition mechanism is ablated, the system can only complete short-term tasks that require fewer steps and shows a reduction in efficiency.
-In contrast, utilizing our comprehensive skill library enables efficient completion of all tasks.
-Table 1: Completion time and success rate (SR) comparison with and without the Recursive Task Decomposition Mechanism (RTDM). All results are shown with the mean and standard deviation over 10 trials.
-Task with RTDM w/o RTDM
-Time(min) SR Time(min) SR
-[Uncaptioned image] $0.3\pm 0.2$ $100\%$ $2.8\pm 2.5$ $100\%$
-[Uncaptioned image] $0.6\pm 0.3$ $100\%$ $4.6\pm 2.5$ $80\%$
-[Uncaptioned image] $1.4\pm 0.5$ $100\%$ $N/A$ $0\%$
-[Uncaptioned image] $4.7\pm 1.3$ $100\%$ $N/A$ $0\%$
-[Uncaptioned image] $6.2\pm 1.6$ $100\%$ $N/A$ $0\%$
-We then evaluate our framework’s performance on eight more complicated resource collection tasks.
-Table 2 compares the completion times between our multi-agent system (3 agents) and single-agent baseline, demonstrating the efficiency gains through coordination.
-Table 2: Completion time (in minutes) of resource collection tasks comparison between multi-agent(MA) and single-agent(SA) systems.
-All results are shown with the mean and standard deviation over 10 trials.
-For detailed task definitions, please refer to Appendix C.
-Task MA Time SA Time
+The game server operates continuously without pausing during interactions with LLMs, so all agents perform in real time.
+See Appendix for more details.
+$3.$ Main Results
+3.3.1. Resource Collection Task
+We evaluate our framework’s performance on eight composite resource collection tasks in Minecraft.
+Each task requires agents to collect multiple items, where each item involves a long chain of prerequisite dependencies.
+For instance, crafting Diamond Armor necessitates first gathering wood, stone, and iron to produce the required tools, and all of which can be automatically accomplished through the recursive decomposition of our skill library.
+Agents only need to collaborate at the skill level.
+Table 1 compares the completion times between our multi-agent system and the single-agent baseline, demonstrating the efficiency gains achieved through coordination.
+Table 1. Average completion time and standard deviation of resource collection tasks comparing multi-agent (MA, fixed at 3 agents) and single-agent (SA) systems over 10 trials.
+Task MA Time (min) SA Time (min)
 Iron Tool Set [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] $7.8\pm 2.1$ $8.5\pm 3.7$
@@
 Storage System [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $10.0\pm 8.9$ $16.7\pm 7.8$
-The experimental results demonstrate that our Minecraft comprehensive skill library, when employed within a multi-agent framework, can efficiently complete various resource collection tasks as well.
+The experimental results demonstrate that our Minecraft comprehensive skill library, when employed within a multi-agent framework, can efficiently complete various resource collection tasks.
 As shown in the comparisons, the system with three agents significantly reduces task completion times compared to the single-agent baseline across most tasks, validating the effectiveness of our approach and showcasing the efficiency benefits of multi-agent collaboration.
-However, the observed non-linear scaling (multi-agent time > single-agent time/3) does not always achieve ideal linear speedup due to two inherent constraints:
-  • Resource dependency chains: Sequential prerequisites in crafting workflows (e.g., iron tools required for diamond mining)
-  • Spatial contention: Overlapping access to resources in close proximity (e.g., multiple agents competing to gather items from the same ore cluster)
-3.3.2 Boss Combat Task
-We conduct comprehensive evaluations across three challenging combat scenarios with varying agent team sizes.
-Experiments are conducted in three predefined scenarios, each featuring an extremely powerful and representative boss from Minecraft’s major dimensions: the Elder Guardian (Overworld), the Wither (Nether), and the Ender Dragon (End).
-The initial inventory of agents remained consistent across all scenarios.
-Their equipment are validated to be collectible within a short time during the resource collection task, and other consumable resources can also be efficiently gathered ultilizing our skill library.
-For further details, please refer to Appendix C.
+However, the performance gain does not scale linearly with the number of agents, which mainly stems from sequential dependency chains and spatial contention.
+3.3.2. Boss Combat Task
+We conduct comprehensive evaluations across three challenging combat scenarios, each involving different agent team sizes and featuring an extremely powerful and representative boss in Minecraft: Elder Guardian, Wither, and Ender Dragon.
+Each task requires agents to dynamically adjust their strategies based on environmental changes.
+For instance, when fighting the Wither, agents must monitor whether it has entered its final phase where it becomes immune to ranged attacks; when battling the Ender Dragon, agents need to first destroy the end crystals that continuously restore its health.
+Prior to each combat scenario, we equip each agent with identical initial supplies, including weapons, armor, and consumables.
+Our proposed comprehensive skill library supports the collection and crafting of these supplies, for example, the crafting of Diamond Armor has already been validated in previous experiments.
+We also observe that although each agent is homogeneous at initialization, heterogeneous division of labor can emerge through multiple rounds of actions and communication within the multi-agent system.
 The performance of the multi-agent system in the boss combat task is summarized in Table 2, demonstrating that our framework achieves high success rates in completing all challenging boss combat tasks with various agent team sizes.
-Table 3: Boss combat performance across three task scenarios.
-All results are shown with the mean and standard deviation over 12 trials.
-For detailed evaluation metrics definitions, please refer to Appendix C.
-Task Scenario # Agents Time(min) Health Ratio Progress SR
+Table 2.
+Performance of boss combat across three task scenarios, reporting mean values and standard deviations of multiple evaluation metrics over 12 trials.
+”#Agents” refers to the number of agents.
+”Time” refers to the minutes taken to complete the combat and achieve victory and is calculated only for successful trials.
+”Health Ratio” refers to the ratio of the remaining health value of the team to its full health.
+’Progress’ refers to the percentage of damage dealt to the boss to its full health.
+See Appendix for detailed metric definitions.
+Task Scenario # Agents Time (min) Health Ratio Progress Success Rate
 Elder Guardian (Overworld) [Uncaptioned image] 3 $2.4\pm 2.1$ $49.8\pm 32.8\%$ $91.4\pm 20.1\%$ $83.3\%$
-5 $1.2\pm 0.8$ $84.4\pm 7.9\%$ $100.0\pm 0.0\%$ $100.0\%$
-10 $1.2\pm 0.3$ $86.9\pm 8.4\%$ $100.0\pm 0.0\%$ $100.0\%$
+5 $1.2\pm 0.8$ $84.4\pm 7.9\%$ $\pm 0.0\% 100.0100.0\%$
+10 $1.2\pm 0.3$ $86.9\pm 8.4\%$ $\pm 0.0\% 100.0100.0\%$
+20 $1.1\pm 0.4$ $89.3\pm 13.4\%$ $\pm 0.0\% 100.0100.0\%$
 Wither (the Nether) [Uncaptioned image] 3 $1.8\pm 0.8$ $18.6\pm 23.4\%$ $71.8\pm 35.9\%$ $41.7\%$
 5 $1.5\pm 0.5$ $53.4\pm 32.7\%$ $88.1\pm 20.6\%$ $75.0\%$
-10 $1.4\pm 0.3$ $69.5\pm 19.0\%$ $100.0\pm 0.0\%$ $100.0\%$
+10 $1.4\pm 0.3$ $69.5\pm 19.0\%$ $\pm 0.0\% 100.0100.0\%$
+20 $1.0\pm 0.1$ $73.4\pm 3.9\%$ $\pm 0.0\% 100.0100.0\%$
 Ender Dragon (the End) [Uncaptioned image] 3 $N/A$ $0.0\pm 0.0\%$ $20.4\pm 18.5\%$ $0.0\%$
@@
 10 $5.2\pm 1.5$ $43.9\pm 24.4\%$ $98.2\pm 5.9\%$ $91.7\%$
-3.3.3 Adversarial PVP Task
-In the adversarial player-vs-player (PVP) task, we conduct a direct comparison between the parallelized and serialized frameworks across various team sizes.
+20 $2.9\pm 0.8$ $67.9\pm 13.6\%$ $\pm 0.0\% 100.0100.0\%$
+3.3.3. Adversarial PVP Task
+Similar to the Boss Combat Task, in Adversarial PVP Task, agents were provided with initial combat resources, and then the two teams of agents engaged in battles with each other.
+We conduct a direct comparison experiment between the parallelized and serialized frameworks across various team sizes.
+Results are recorded when one team of agents is entirely defeated.
+After multiple experiments, victory rates and other relevant metrics are calculated.
+To promote fair confrontation, we provide each team with exactly the same initial supplies.
 As shown in Fig.
 3, this setup clearly demonstrate and quantify the differences of these two methodologies under competitive conditions.
-As shown in Fig.
-3, the parallelized framework demonstrates a significant advantage over the serialized framework in this dynamic adversarial scenario.
-Our analysis indicates that this advantage is primarily attributed to our interruption mechanism, which enables agents to dynamically adjust their strategies and respond promptly to changes in the environment (e.g., seamlessly switch attack targets, prioritize health restoration, etc.).
-Figure 3: Comparison of parallelized vs. serialized frameworks in PVP tasks. All results are shown with the mean and standard deviation over 10 trials.
-3.4 Ablation Study
+The parallelized framework demonstrates a significant advantage over the serialized framework in this dynamic adversarial scenario.
+Our analysis indicates that this advantage is primarily attributed to our interruption mechanism, which enables agents to dynamically adjust their strategies and respond promptly to changes in the environment (e.g., seamlessly switch attack targets, prioritize health restoration, more efficient communication, etc.).
+Figure 3. Comparison of parallelized vs. serialized frameworks in PVP tasks. All results are shown with the mean and standard deviation over 10 trials.
+3.4. Ablation Study
+3.4.1. Recursive Task Decomposition
+We first evaluate the effectiveness of our skill library’s recursive task decomposition mechanism by assessing a single agent’s performance on resource collection tasks in Minecraft.
+This evaluation aimed to verify the comprehensive skill library’s capability in automating complex workflows, thereby validating its efficiency and reliability in handling fundamental Minecraft tasks.
+As shown in Table.
+3, when the recursive task decomposition mechanism is ablated, the system can only complete short-term tasks that require fewer steps and shows a reduction in efficiency.
+In contrast, utilizing our comprehensive skill library with the recursive task decomposition mechanism enables efficient completion of all tasks.
+Table 3.
+Average completion time with standard deviation and success rate (SR) comparison with and without the recursive task decomposition mechanism (RTDM).
+Tasks from top to bottom: (1) crafting table, (2) wooden tool, (3) stone tool, (4) iron tool, (5) diamond.
+Task with RTDM w/o RTDM
+Time (min) SR Time (min) SR
+[Uncaptioned image] $0.3\pm 0.2$ $100\%$ $2.8\pm 2.5$ $100\%$
+[Uncaptioned image] $0.6\pm 0.3$ $100\%$ $4.6\pm 2.5$ $80\%$
+[Uncaptioned image] $1.4\pm 0.5$ $100\%$ $N/A$ $0\%$
+[Uncaptioned image] $4.7\pm 1.3$ $100\%$ $N/A$ $0\%$
+[Uncaptioned image] $6.2\pm 1.6$ $100\%$ $N/A$ $0\%$
+In the more complex resource collection task we propose in Section 3.3.1, the MAS will not be able to make any meaningful progress without the recursive task decomposition mechanism, suggesting that invoking an LLM to independently plan the task dependencies embedded within a DAG would not be as efficient as utilizing a fixed mechanism.
+3.4.2. Parallelized Framework
 To validate the necessity of our method, we perform ablation studies in boss combat tasks by disabling different components in our framework:
-  • w/o Parallelized Planning-Acting Framework: Replaced our concurrent architecture with traditional serialized execution (LLM call → action execution), disabling action interruption capabilities once execution begins.
-  • w/o Centralized Memory System: Disabled real-time team observation polling, team chat logs, action history and global progress information, thereby restricting agents to rely solely on their individual observations.
-  • Multimodal Observation: Replaced text-based observations with visual inputs using Mineflayer’s prismarine-viewer, capturing first-person screenshots processed by a vision-language model (VLM) as observation.
-These ablation configurations systematically evaluate the contribution of each architectural component to overall system performance, providing insights into their relative importance in complex multi-agent coordination tasks.
+  • w/o Parallelized Planning-Acting Framework (PPA): Replaced our parallelized framework with traditional serialized execution, disabling the interruption mechanism.
+  • w/o Centralized Memory System (CMS): Disabled real-time team observation polling, chat logs, action history and global progress information, thereby restricting agents to rely solely on their individual observations.
 Experiment results shown in Fig.
-4 highlight the critical roles of the parallelized planning-acting framework and centralized memory system in maintaining this high performance.
-When we replaced the text-based observation space with a multimodal observation system, we observed that although the inclusion of visual language models (VLMs) increased response latency and lowered the observation accuracy, the agents still performed reasonably well.
-This indicates that our framework remains robust even when handling different modalities.
-Figure 4: Ablation study in boss combat tasks across 12 trials in three scenarios: Elder Guardian (EG), Wither (Wi), and Ender Dragon (ED), where efficiency is defined as the inverse of completion time in minutes.
-3.5 Human-Agent Interaction
-Our framework also supports flexible human-agent interaction, allowing human players to either actively participate in task execution alongside agents, facilitating dynamic cooperation within the game environment, or to take on a guiding role by providing instructions to agents for task division, coordination, and strategic decision-making.
-Further experimental cases demonstrating the effectiveness of human-agent interaction are presented in Appendix C.4.
-4 Related Works
+4 highlight the critical roles of the parallelized planning-acting framework and the centralized memory system.
+Besides these ablation studies, our method also performs exceptionally well in direct comparative experiments.
+Please refer to Appendix for details.
+Figure 4.
+Ablation study in boss combat tasks across 12 trials in three scenarios: Elder Guardian (EG), Wither (Wi), and Ender Dragon (ED), where efficiency is defined as the inverse of completion time in minutes.
+Table 4.
+Performance of boss combat across three task scenarios using different observation modalities, with a fixed team size of 10 agents.
+Reporting mean values and standard deviations of multiple evaluation metrics over 12 trials.
+Task Scenario Observation Modality Time(min) Health Ratio Progress Success Rate
+Elder Guardian Text $1.2\pm 0.3$ $86.9\pm 8.4\%$ $\pm 0.0\% 100.0100.0\%$
+Visual $1.9\pm 0.7$ $85.3\pm 6.1\%$ $\pm 0.0\% 100.0100.0\%$
+Wither Text $1.4\pm 0.3$ $69.5\pm 19.0\%$ $\pm 0.0\% 100.0100.0\%$
+Visual $2.3\pm 0.5$ $39.1\pm 12.4\%$ $82.8\pm 7.2\%$ $66.7\%$
+Ender Dragon Text $5.2\pm 1.5$ $43.9\pm 24.4\%$ $98.2\pm 5.9\%$ $91.7\%$
+Visual $6.1\pm 1.1$ $21.8\pm 10.3\%$ $58.3\pm 13.5\%$ $50.0\%$
+3.5. Robustness Analysis
+3.5.1. Robust to Modality
+We evaluated the framework’s robustness to different modalities by replacing the text-based observation with a multi-modal approach in boss combat tasks.
+While the inclusion of visual language models (VLMs) introduced increased response latency and slightly reduced observation accuracy, the agents maintained reasonable performance levels.
+As summarized in Table 4, our framework achieves strong performance across all three challenging task scenarios using either modalities.
+3.5.2. Robust to Scale
+To validate the effectiveness of our approach under large-scale scenarios, we conducted additional experiments involving teams of 5-50 agents.
+These experiments focused on measuring the impact of agent quantity (where the number of agents equaled the number of latest chat entries read per LLM inference) on token counts and inference time.
+Figure 5. Inference time and input/output tokens per single-agent planning step of MAS at different scales.
+In Fig.
+5, it can be observed that the inference time of the LLM tends to stabilize instead of increasing continuously as the number of agents increases.
+Additionally, the total token cost shows an approximately linear growth with the number of agents, which is both acceptable and manageable.
+Since each agent is equipped with parallelized planning threads and acting threads independently, the actions of each agent remain unaffected by the total number of agents in the system.
+Further analysis can be found in the Appendix.
+The input tokens increases linearly with the number of agents, as more chat entries are processed during each planning step.
+However, the output tokens remain stable due to the constraints imposed by the prompt format, which limits the length of generated responses.
+Overall, the system’s total latency does not increase significantly with the number of agents, and the total cost (tokens) grows approximately linearly with the number of agents, which is acceptable and controllable.
+4. Related Works
 LLM-based Minecraft Agents.
@@
 (2023) established the first LLM-based agent with automatic skill discovery using GPT-4 Achiam et al.
-(2023).
+(2023), built on the open-source library Mineflayer PrismarineJS (2023).
 Subsequent studies enhanced agents via specialized memory mechanisms Park et al.
@@
 (2023); Cai et al.
-(2024), along with other novel techniques  Wang et al.
+(2024), along with other novel techniques Li et al.
+(2025); Wang et al.
 (2024); Zhou et al.
@@
 (2024); Dong et al.
+(2024); Wei et al.
 (2025).
 LLM-based Multi-Agent Systems.
-Recent surveys  Mou et al.
+Recent research has focused on several core areas Mou et al.
 (2024); Guo et al.
-(2024) provide a comprehensive overview of the key advancements in LLM-based multi-agent systems.
-Building on these foundations, recent research has focused on several core areas: infrastructure frameworks for efficient agent coordination  Gong et al.
+(2024): infrastructure frameworks for efficient agent coordination Hong et al.
 (2023); Chen et al.
-(2024b); Zhang et al.
-(2024b); Hong et al.
+(2024c); Zhang et al.
+(2024a, b, 2025a), which introduce novel paradigms for task management and team collaboration; efforts aimed at improving MAS collaboration efficiency through optimizations at communication or routing Yue et al.
 (2025); Zhang et al.
-(2024a), which introduce novel paradigms for task management and team collaboration; benchmark development to evaluate multi-agent performance in dynamic environments  Chen et al.
-(2024a); Dong et al.
+(2025b); Chen et al.
+(2025); benchmark development to evaluate multi-agent performance in dynamic environments Chen et al.
+(2024b); Dong et al.
 (2024), which has created robust testing environments to assess the generalization and efficiency of LLM-powered agents; large-scale social simulations AL et al.
@@
 (2024); Chen et al.
-(2024c) demonstrating the effectiveness of LLM-based agents in specific scenario simulation.
-5 Conclusion
-We propose a novel parallelized planning-acting multi-agent framework that significantly enhances the responsiveness and adaptability of LLM-based MAS in dynamic environments.
+(2024a) demonstrating the effectiveness of LLM-based agents in specific scenario simulation.
+In contrast to prior MAS research that primarily focuses on coordination paradigms or communication strategies, our work targets continuous real-time interaction in non-paused dynamic environments.
+5. Conclusion
+We propose a novel parallelized planning-acting multi-agent framework that significantly enhances the responsiveness and adaptability of LLM-based MAS in dynamic environments like Minecraft.
 Our framework’s dual-thread architecture with interruptible execution mechanism enables real-time interaction and continuous adaptation, overcoming the limitations of traditional serialized execution paradigms.
-The comprehensive skill library and automated task decomposition mechanism further improve efficiency and coordination as an engineering contribution.
-Through extensive experiments, our framework demonstrates superior performance in both collaborative and adversarial scenarios, and achieves more immediate human-agent interaction, highlighting its potential for broader applications.
-6 Limitations
-While demonstrating significant advantages, our framework has three main limitations.
-First, the continuous LLM request design leads to higher computational costs, which may limit scalability in resource-constrained environments.
-Second, LLM hallucinations occasionally result in irrational action interruptions, affecting the reliability of agent behavior.
-Third, although our framework adapts to multimodal observations, the current fusion method relies solely on visual-language models (VLMs), potentially failing to fully exploit the complementarity of multimodal information and limiting performance in complex scenarios.
-For future work, we aim to optimize computational efficiency by refining LLM request mechanisms, mitigate hallucination effects through enhanced verification strategies, and improve multimodal fusion techniques to better leverage diverse sensory inputs.
+The comprehensive skill library and recursive task decomposition mechanism further improve efficiency and coordination as an engineering contribution.
+Experiments on our challenging benchmark tasks validate the effectiveness of our framework in diverse Minecraft scenarios.
+{acks}
+This work is supported in part by the Hangzhou Joint Funds of the Zhejiang Provincial Natural Science Foundation of China under Grant No.
+LHZSD24F020001, in part by the Zhejiang Province High-Level Talents Special Support Program “Leading Talent of Technological Innovation of Ten-Thousands Talents Program” under Grant No.
+2022R52046, in part by the Fundamental Research Funds for the Central Universities under Grant No.
+2021FZZX001-23, and in part by the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.
 Appendix A Centralized Memory System
-The memory of the multi-agent system is stored in JSON files, and agents read and modify their contents by reading and writing to these files. Below is an example of how the memory is stored:
-A.1 Agent Observation & Status
-"time": "day",
-"health": 12.8,
-"hunger": 17,
-"position": "(10, 65, -8)",
-"equipment": [
-null,
-null,
-null,
-null,
-"bow",
-null
-"inventory": {
-"arrow": 187,
-"golden_apple": 4,
-"diamond_pickaxe": 1,
-"diamond_sword": 1,
-"diamond_helmet": 1,
-"diamond_chestplate": 1,
-"diamond_leggings": 1,
-"diamond_boots": 1,
-"bow": 1
-"nearbyBlocks": [
-"end_stone",
-"bedrock"
-"nearbyEntities": "enderman, ender_dragon",
-"nearbyPlayers": [
-"Steve",
-"Brine",
-"Jeb",
-"Notch"
-A.2 Team Chat Log
-"Day 1 - Morning": [
-"SystemInfo: Number of End Crystals Remaining: 10",
-"SystemInfo: Boss Health: 200",
-"SystemInfo: Boss Health: 200",
-"Notch: Lets prioritize taking down the end_crystals to prevent the Ender Dragon from healing.
+The memory of the multi-agent system is stored in JSON files, and agents read and modify their contents by reading and writing to these files.
+The observations and states of the agent are updated in a polling manner per second and saved in the following form, constantly overwritten.
+Observation Records { time”: day”, health”: 12.8, hunger”: 17, position”: ”(10, 65, -8)”, equipment”: [ null, null, null, null, bow”, null ], inventory”: { arrow”: 187, golden_apple”: 4, diamond_pickaxe”: 1, diamond_sword”: 1, diamond_helmet”: 1, diamond_chestplate”: 1, diamond_leggings”: 1, diamond_boots”: 1, bow”: 1 }, nearbyBlocks”: [ end_stone”, bedrock ], nearbyEntities”: enderman, ender_dragon”, nearbyPlayers”: [ Steve”, Brine”, Jeb”, Notch ] }
+The team chat logs are constantly updated by active communication and continue to grow in real time, but only the most recent information is taken as prompts.
+Chat Logs (Active Communication) { Day 1 - Morning”: [ SystemInfo: Number of End Crystals Remaining: 10”, SystemInfo: Boss Health: 200”, SystemInfo: Boss Health: 200”, Notch: Lets prioritize taking down the end_crystals to prevent the Ender Dragon from healing.
 I will start attacking the nearest end_crystal with my bow.
-Everyone, cover me and be ready to assist if needed.",
-"Steve: Im ready with my bow and arrows.
+Everyone, cover me and be ready to assist if needed.”, Steve: Im ready with my bow and arrows.
 Lets focus on taking down the end_crystals first to prevent the Ender Dragon from healing.
 Ill start targeting the nearest end_crystal.
-Stay alert for any incoming endermen but avoid engaging them unless necessary.",
-"SystemInfo: Boss Health: 200",
-"Brine: Brine: Lets prioritize destroying the end_crystals to prevent the Ender Dragon from healing.
+Stay alert for any incoming endermen but avoid engaging them unless necessary.”, SystemInfo: Boss Health: 200”, Brine: Brine: Lets prioritize destroying the end_crystals to prevent the Ender Dragon from healing.
 Notch and I will start targeting the nearest crystal.
@@
 Stay alert for any Endermen but avoid engaging them unless necessary.
-Report if you need help or spot any crystals.",
-"Jeb: Im ready with my bow and arrows.
+Report if you need help or spot any crystals.”, Jeb: Im ready with my bow and arrows.
 Lets prioritize the end_crystal on the obsidian pillars to prevent the Ender Dragon from healing.
 Ill start targeting the nearest crystal.
-Stay alert for any incoming endermen but avoid engaging unless necessary."
-A.3 Team Observation Log
-"obs_entries": [
-"Notch: Team, Im fully equipped with diamond armor, a bow, and 200 arrows.
+Stay alert for any incoming endermen but avoid engaging unless necessary.” ] }
+The team chat logs are also continuously updated in real time by passive communication, but only the latest observations of each agent are provided to the LLM as prompts.
+Chat Logs (Passive Communication) { obs_entries”: [ Notch: Team, Im fully equipped with diamond armor, a bow, and 200 arrows.
 I see an end_crystal and an ender_dragon nearby, along with some endermen.
@@
 Everyone, prepare your bows and stay alert for any incoming endermen.
-Ready when you are!",
-"Brine: Team, Im currently well-equipped with a bow, 200 arrows, and diamond armor.
+Ready when you are!”, Brine: Team, Im currently well-equipped with a bow, 200 arrows, and diamond armor.
 The ender_dragon is nearby but hasnt attacked yet.
@@
 Stay alert for any Endermen, but avoid engaging them unless necessary.
-Report if you spot any crystals or need assistance.",
-"Jeb: All end_crystals have been destroyed.
+Report if you spot any crystals or need assistance.”, Jeb: All end_crystals have been destroyed.
 Focusing on the Ender Dragon now.
@@
 Stay alert for any incoming endermen but prioritize the Ender Dragon.
-Lets finish this!",
-"Steve: Im in critical condition and need immediate assistance!
-My health is down to 0.
-Brine, Notch, Jeb, please cover me or help me get to safety.
-I have a full bow and golden apples but cant fight effectively in this state.
-Endermen are nearby, be cautious!",
-A.4 Team Action Log
-"Notch": [
+Lets finish this!”, ] }
+The team action logs record the historical actions and reasoning processes of each agent, but only an agent’s own current action is used as LLM prompts.
 […]
-[…]
-[…]
-"Steve": ...
-Part of above information is provided as human prompts for agents to plan actions.
 In addition to this, agents are also provided with identity information such as their name, team affiliation, and teammates’ names, to simulate a human-like multi-agent collaborative scenario.
-Appendix B Minecraft Comprehensive Skill Library
-Our proposed skill library is implemented based on MineflayerPrismarineJS (2023).To simulate the process of solving prerequisite steps in the actual game to eventually obtain specific items, this skill library has introduced a recursive task decomposition mechanism.
+Appendix B Comprehensive Skill Library
+Our proposed comprehensive skill library for Minecraft is implemented based on Mineflayer.
+To simulate the process of solving prerequisite steps in the actual game to eventually obtain specific items, this skill library has introduced a recursive task decomposition mechanism.
 From the implementation level, all interfaces can recursively invoke each other or themselves and pass parameters.
-B.1 Basic Skill Interfaces
-  • obtainItem(bot, count, type): Automates the collection of items.
+B.1. Minecraft Skill APIs
+The Basic Skill Interfaces provide the foundational functions required to perform essential actions in the game, such as collecting items, crafting, mining, smelting, and combat.
+These functions are designed to be modular and reusable, allowing for seamless integration into more complex workflows.
+Basic Skill Interfaces obtainItem(bot, count, type): Automates the collection of items.
 mineItem(bot, count, type, explore_direction, explore_time): Mines using specific tools.
@@
 The exploration time limit defaults to five minutes (6000 ticks in Minecraft).
-  • craftItem(bot, count, type, need_crafting_table): Crafts items. If no parameter is specified, the crafting is done using a crafting table by default.
-  • smeltItem(bot, count, type, fuel): Smelts or cooks items in a furnace. If no parameters are provided, coal is used as the default fuel.
+craftItem(bot, count, type, need_crafting_table): Crafts items.
+If no parameter is specified, the crafting is done using a crafting table by default.
+smeltItem(bot, count, type, fuel): Smelts or cooks items in a furnace.
+If no parameters are provided, coal is used as the default fuel.
 collectItem(bot, count, type, function): Collects items by killing animals or using other special methods.
@@
 listenChat(bot, player_name): Continuously listens to a player’s chat messages.
-B.2 Recursive-related Interfaces
-  • preTool(item): Retrieves the minimum prerequisite tools required to gather a specific item.
+The Recursive-related Interfaces are designed to handle the dependencies between tasks.
+These functions determine the prerequisites for obtaining specific items, such as tools, materials, or entities, and ensure that all necessary steps are completed before proceeding with the main task.
+All task dependencies are modeled as a Directed Acyclic Graph (DAG), where nodes represent tasks and edges represent prerequisite relationships.
+We have cataloged thousands of these intrinsic dependencies within Minecraft to support this recursive decomposition mechanism.
+This not only provides a methodological framework but also constitutes a significant engineering contribution by facilitating the automation of complex task sequences.
+Recursive-related Interfaces preTool(item): Retrieves the minimum prerequisite tool required for a specific item.
 preItem(item): Retrieves the prerequisite items required to craft a specific item and indicates whether a crafting table is necessary.
 preSmelt(item): Retrieves the prerequisite items required to smelt a specific item.
-  • preCollect(item): Retrieves the prerequisite entities required to collect a specific item.
+preCollect(item): Retrieves the prerequisite entities required for a specific item.
 getFunc(item): Retrieves the method used to collect a specific item.
-This library is scalable and can be easily extended to accommodate new items, recipes, and tasks.
-Currently, it supports the collection of over 790 types of items across all three dimensions of Minecraft (Overworld, Nether, and End).
-However, the library does not yet support some specialized gameplay features.
-Future updates will aim to expand the library’s functionality, adding support for more complex tasks.
+B.2. Implementation Details
+Algorithm 2 demonstrates the implementation of the obtainItem function, which determines the collection method for a specific item and delegates the task to the corresponding sub-function (e.g., mining, crafting, smelting).
+This function serves as the entry point for recursive task decomposition.
+Algorithm 2 obtainItem
+1:function obtainItem($bot,cnt,type$)
+2:  $func\leftarrow\textsc{getCollectionMethod}(type)$
+3:  if $func=\text{``craft''}$ then
+4:   return $\textsc{craftItem}(bot,cnt,type)$
+5:  else if $func=\text{``mine''}$ then
+6:   return $\textsc{mineItem}(bot,cnt,type)$
+7:  else if $func=\text{``smelt''}$ then
+8:   return $\textsc{smeltItem}(bot,cnt,type)$
+9:  else
+10:   return false
+11:  end if
+12:end function
+The mineItem function shown in Algorithm 3 handles the process of mining blocks or ores using appropriate tools.
+If the required tool is not in the inventory, it recursively calls obtainItem to craft or gather the tool before proceeding with mining.
+This ensures that the bot is always equipped with the necessary tools for the task.
+Algorithm 3 mineItem
+1:function mineItem($bot,cnt,type,dir$)
+2:  $tool\leftarrow\textsc{preTool}(type)$
+3:  if $\textsc{invCnt}(tool)=0$ then
+4:   $\textsc{obtainItem}(bot,1,tool)$
+5:  end if
+6:  $\textsc{equip}(bot,tool)$
+7:  if $dir=\text{null}$ then
+8:   $dir\leftarrow\textsc{randomDirection}$
+9:  end if
+10:  $\textsc{startMining}(bot,type,dir,time)$
+11:  return $\textsc{verifyResult}(cnt,type)$
+12:end function
+The craftItem function shown in Algorithm 4 retrieves the recipe requirements for the specified item.
+If the required materials are insufficient, it recursively calls obtainItem to gather them before proceeding with crafting.
+Algorithm 4 craftItem
+1:function craftItem($bot,cnt,type$)
+2:  $reqs\leftarrow\textsc{getRecipeReq}(type)$
+3:  for (r_cnt,r_type)(r\_cnt,r\_type) in reqsreqs do
+4:   if $\textsc{invCnt}(r\_type)<r\_cnt$ then
+5:     $n\leftarrow r\_cnt-\textsc{invCnt}(r\_type)$
+6:     $\textsc{obtainItem}(bot,n,r\_type)$
+7:   end if
+8:  end for
+9:  $\textsc{executeCrafting}(type,cnt)$
+10:  return $\textsc{verifyResult}(cnt,type)$
+11:end function
+The smeltItem function shown in Algorithm 5 checks if input items for smelting are available. If not, it recursively calls obtainItem to gather necessary materials before proceeding with smelting.
+Algorithm 5 smeltItem
+1:function smeltItem($bot,cnt,type$)
+2:  $input\leftarrow\textsc{getSmeltingInput}(type)$
+3:  if $\textsc{invCnt}(input)<cnt$ then
+4:   $n\leftarrow cnt-\textsc{invCnt}(input)$
+5:   $\textsc{obtainItem}(bot,n,input)$
+6:  end if
+7:  prepareFurnaceAndFuel
+8:  $\textsc{executeSmelting}(type,cnt)$
+9:  return $\textsc{verifyResult}(cnt,type)$
+10:end function
 Appendix C Experiment Details
-Experiments were conducted using the Qwen-Plus model API provided by Alibaba Cloud.
-In the multiplayer local area network (LAN) server of Minecraft, agents connected to the game via different ports and interacted with the environment and other players.
-C.1 Resource Collection Task
-We implemented a comprehensive Minecraft resource collection benchmark, which allows multi-agent collaboration to automatically collect resources based on a given requirement dictionary, making this benchmark general and scalable.
-We conducted experiments on the following representative resource collection tasks:
-Task Definitions:
-  • Iron Tool Set: { ’iron_pickaxe’: 1, ’iron_shovel’: 1, ’iron_hoe’: 1, ’iron_axe’: 1 } — A set of commonly used iron tools in Minecraft.
+All experiments were conducted using the Qwen-Plus and Qwen-VL-plus model API provided by Alibaba Cloud.
+In the multiplayer local area network (LAN) server of Minecraft Java Edition version 1.19.4, agents connected to the game via different ports as an independent player and interacted with the environment and other agent players.
+C.1. Resource Collection Task
+Our framework supports tasks formulated as collecting any quantity of various types of items, represented by a requirement dictionary.
+This design ensures the richness and flexibility of supported tasks, enabling multi-agent collaboration to automatically collect resources based on given requirements.
+Such a setup makes our benchmark both general and scalable.
+However, due to inevitable time and API costs, we selected a set of representative resource collection tasks for our experiments.
+These tasks not only demonstrate the capabilities of our framework but also highlight common challenges faced by players in Minecraft, such as gathering essential survival tools (the Iron Tool Set task) or preparing for combat with powerful boss monsters (the Diamond Armor task).
+The following are the representative resource collection tasks we experimented with:
+Task Definitions Iron Tool Set: { ’iron_pickaxe’: 1, ’iron_shovel’: 1, ’iron_hoe’: 1, ’iron_axe’: 1 } — A set of commonly used iron tools in Minecraft.
 Diamond Armor: { ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1 } — A full set of diamond armor in Minecraft.
@@
 Each round of the experiment is conducted in a world generated with a random seed to ensure the generalizability of the results.
-C.2 Boss Combat Task
-We manually implemented three representative and challenging task scenarios: defeating the Elder Guardian in an ocean monument in the Overworld, defeating the Wither in the Nether, and defeating the Ender Dragon in the End.
-Defeating the Ender Dragon represents the final achievement in the game.
+C.2. Boss Combat Task
+In Minecraft, there are three primary world dimensions: Overworld, Nether, and End.
+Each dimension hosts powerful boss monsters whose defeat is considered the pinnacle challenge for players of Minecraft.
+Based on this, we have predefined three combat tasks:
+  • Elder Guardian (Overworld): This task involves defeating the formidable boss Elder Guardian within complex underwater terrain of an ocean monument.
+Agents must also contend with other surrounding monsters, adding complexity that requires the MAS to develop effective combat strategies.
+  • Wither (Nether): This task involves battling the powerful boss Wither in the Nether.
+Agents face additional challenges from smaller enemies such as piglins and ghasts, amidst a landscape filled with lava and other hazards.
+When the Wither’s health drops below half, it enters a berserk state where it becomes immune to ranged attacks, necessitating the MAS to dynamically adjust combat strategies based on real-time observations.
+  • Ender Dragon (End): Defeating the Ender Dragon in the End is considered the ultimate challenge in Minecraft.
+Multiple agents need to strategically cooperate by first destroying the end crystals located atop obsidian pillars to disable the dragon’s healing mechanism before engaging in battle with the dragon and its surrounding endermen, requiring the MAS to adjust tactics adaptively.
 Prior to each combat, agents were equipped with standardized combat resources, including a full set of diamond armor (validated as efficiently collectible through our resource collection experiments), bow and arrows, and some consumables.
@@
 Task completion time was measured excluding initialization phases.
-Evaluation metrics are defined as follows: health ratio = $\frac{1}{n}\sum_{i=1}^{n}\frac{h_{i}}{H_{\text{max}}}\times 100\%$ ($h_{i}$: agent’s remaining health, $H_{\text{max}}$: agent’s max health) and progress = $\frac{H_{\text{boss}}-h_{\text{boss}}}{H_{\text{boss}}}\times 100\%$ ($h_{\text{boss}}$: boss’s remaining health, $H_{\text{boss}}$: boss’s max health).
-C.3 Adversarial PVP Task
+Evaluation metrics are defined as follows:
+$\text{health ratio}=\frac{1}{n}\sum_{i=1}^{n}\frac{h_{i}}{H_{\text{max}}}\times 100\%$ (10)
+where $h_{i}$ is the remaining health of agent $i$, and $H_{\text{max}}$ is the maximum health of the agent.
+$\text{progress}=\frac{H_{\text{boss}}-h_{\text{boss}}}{H_{\text{boss}}}\times 100\%$ (11)
+where $h_{\text{boss}}$ is the remaining health of the boss monster, and $H_{\text{boss}}$ is the maximum health of the boss monster.
+C.3. Adversarial PVP Task
+In addition to ablation studies, we design a task scenario where two frameworks can be directly compared, namely Adversarial PVP Task.
 Similar to the Boss Combat Task, in this task, agents were provided with combat resources.
+Then, the two teams of agents engaged in battles with each other.
 Results were recorded when one team of agents was entirely defeated.
@@
 Here is an example of initial inventory given to an agent : { ’wooden_sword’: 1, ’diamond_helmet’: 1, ’diamond_chestplate’: 1, ’diamond_leggings’: 1, ’diamond_boots’: 1, ’bow’: 1, ’arrow’: 200, ’golden_apple’: 20 }
-C.4 Human-Agent Interaction
-This section presents examples of human-agent interactions in dynamic environments, highlighting how agents and human players collaborate in various tasks.
-Scenario 1: Human-Agent Collaboration in Resource Collection Tasks
-Agent Steve’s Environment Information:
-"time": "day",
-"health": 20,
-"hunger": 20,
-"position": "(-223, 60, -159)",
-"equipment": [
-null,
-null,
-null,
-null,
-"crafting_table",
-null
-"inventory": {
-"crafting_table": 1
-"nearbyBlocks": [
-"dirt",
-"grass_block",
-"oak_log",
-"oak_leaves"
-"nearbyEntities": "pig",
-"nearbyPlayers": ["Alex", "Notch"]
-Human Player’s Chat Message:
-’Hey Steve, I’ve got some beef, can you hunt a pig for obtaining porkchops?’
-Agent Steve’s Planning Action:
-"skill": "obtainItem(bot, 10, \’porkchop\’)",
-"interrupt": false,
-"reason": "Following the human players request to hunt a pig for obtaining porkchops. Since there are pigs nearby, I will proceed with hunting to gather porkchops."
-Scenario 2: Human Player Directing Agents for Combat Tasks
-Agent Steve’s Environment Information:
-"time": "day",
-"health": 1.2,
-"hunger": 17,
-"position": "(6, 66, 13)",
-"equipment": [
-"diamond_helmet",
-"diamond_chestplate",
-"diamond_leggings",
-"diamond_boots",
-"bow",
-null
-"inventory": {
-"arrow": 189,
-"golden_apple": 4,
-"wooden_sword": 1,
-"bow": 1
-"nearbyBlocks": [
-"end_stone",
-"bedrock"
-"nearbyEntities": "ender_dragon",
-"nearbyPlayers": [
-"Steve",
-"Notch",
-"Jeb"
-Human Player’s Chat Message:
-’Steve, you are in grave danger. Cease your attack immediately and prepare to restore health!’
-Agent Steve’s Planning Action:
-"skill": "consumeItem(bot, \’golden_apple\’, false)",
-"interrupt": true,
-"reason": "My health is critically low at 1.2, and I need to restore it immediately to continue fighting the ender_dragon effectively. I will follow the human players advice and prioritize my health."
+C.4. Ablation Study
+Our ablation studies respectively focused on the comprehensive skill library and its recursive task decomposition, the centralized memory system and the parallelized planning-acting mechanism, which are three core contributions proposed in our paper.
+Although it is challenging to conduct fair comparisons between different approaches and frameworks in agent-based research due to varying architectures and objectives, each ablation here can be regarded as a methodological baseline.
+For example, ablating the parallelized planning-acting mechanism is actually taking the traditional serialized planning-acting architecture as the baseline.
+C.5. Scale Robustness Analysis
+Our method theoretically supports multi-agent systems of any scale.
+Regardless of the scale, each agent is independently equipped with parallel planning threads and acting threads and interacts with the environment in real time.
+As for the acting thread of a single agent, its actions are completely unaffected by the number of agents in the entire system.
+However, for the planning thread, under our settings, the length of the chat log window input for each LLM call is equal to the number of system agents, which means that as the system scale increases, the number of input tokens for LLM inference continues to increase.
+So we conducted an experiment on this.
+By adjusting the window length of the chat log, we sampled different numbers of chat entries for LLM planning and counted their inference duration and the number of input and output tokens.
+As shown in the main paper, experiment results show that the LLM inference time tends to stabilize rather than grow continuously with increasing agents and the total token cost grows approximately linearly with the number of agents.
+This may be attributed to the fact that we utilize the Qwen-Plus API, which likely incorporates advanced optimization techniques such as vLLM acceleration, enabling the system to efficiently handle increased tokens without significantly impacting inference latency, thereby maintaining stable performance even as the input size grows.
+The input tokens increases linearly with the number of agents, as more chat entries are processed during each planning step.
+However, the output tokens remain stable due to the constraints imposed by the prompt format, which limits the length of generated responses.
+Overall, our method demonstrates robustness to the scale of multi-agent systems.
+Appendix D Human-Agent Interaction
+Our framework also supports flexible human-agent interaction, allowing human players to either actively participate in task execution alongside agents, facilitating dynamic cooperation within the game environment, or to take on a guiding role by providing instructions to agents for task division, coordination, and strategic decision-making.
+This section presents examples of human-agent interactions in Minecraft, highlighting how AI agents and human players collaborate in various tasks.
+D.0.1. Scenario 1: Human-Agent Collaboration
+Agent Steve’s Environment Information { time”: day”, health”: 20, hunger”: 20, position”: ”(-223, 60, -159)”, equipment”: [ null, null, null, null, crafting_table”, null ], inventory”: { crafting_table”: 1 }, nearbyBlocks”: [ dirt”, grass_block”, oak_log”, oak_leaves ], nearbyEntities”: pig”, nearbyPlayers”: [”Alex”, Notch”] }
+Human Player’s Chat Message ’Hey Steve, I’ve got some beef, can you hunt a pig for obtaining porkchops?’
+Agent Steve’s Planning Action { skill”: obtainItem(bot, 10, \’porkchop\’)”, interrupt”: false, reason”: Following the human players request to hunt a pig for obtaining porkchops.
+Since there are pigs nearby, I will proceed with hunting to gather porkchops.” }
+Scenario 2: Human Player Directing Agents
+Agent Steve’s Environment Information { time”: day”, health”: 1.2, hunger”: 17, position”: ”(6, 66, 13)”, equipment”: [ diamond_helmet”, diamond_chestplate”, diamond_leggings”, diamond_boots”, bow”, null ], inventory”: { arrow”: 189, golden_apple”: 4, bow”: 1 }, nearbyBlocks”: [ end_stone ], nearbyEntities”: ender_dragon”, nearbyPlayers”: [ Jeb ] }
+Human Player’s Chat Message ’Steve, you’re in grave danger. Cease attack immediately and prepare to restore health!’
+Agent Steve’s Planning Action { skill”: consumeItem(bot, \’golden_apple\’, false)”, interrupt”: true, reason”: My health is critically low at 1.2, and I need to restore it immediately to continue fighting the ender_dragon effectively.
+I will follow the human players advice and prioritize my health.” }